### v1

# The Multi-Wavelength Context of Delayed Radio Emission in TDEs: Evidence for Accretion-Driven Outflows

Abstract
Recent observations presented in Cendes et al. (2024a) show that optically selected tidal disruption events (TDEs) commonly produce delayed radio emission that can peak years post-disruption. Here, we explore the multi-wavelength properties of a sample of radio-observed optically selected TDEs to shed light on the physical process(es) responsible for the late-rising radio emission. We combine new late-time X-ray observations with archival optical, UV, X-ray, and radio data to conclude that a diversity of accretion-driven outflows may power delayed radio emission in TDEs. Our analysis suggests that some late radio outflows may be launched by a delayed phase of super-Eddington accretion onto the central supermassive black hole (SMBH), while others may result from a state transition to a “low-hard” radiatively inefficient accretion flow or the deceleration of an off-axis relativistic jet. We additionally find weak statistical evidence that TDEs with delayed radio emission have larger optical/UV photospheric radii than other TDEs and are less likely to exhibit helium emission lines at early times, possibly also supporting the hypothesis that the onset of SMBH accretion is delayed in these systems. Our results have implications for our understanding of state changes in SMBH accretion flows, the circularization timescale for TDE debris, and the prevalence of off-axis jets in TDEs, and motivates systematic, long-term monitoring of these unique transients. The brightest objects in our sample are also detected in the VLA Sky Survey (VLASS), demonstrating that all-sky radio surveys can play an important role in discovering unexpected radio properties of the TDE population.
facilities: VLA, MeerKAT, ATCA, Swift(XRT and UVOT), Chandra, XMM-Newtonsoftware: astropy (Astropy Collaboration et al., 2013, 2018), mosfit (Guillochon et al., 2018), Source Extractor (Bertin & Arnouts, 1996)
1 Introduction
The strong gravitational field near a supermassive black hole (SMBH) is capable of tidally disrupting an approaching star, temporarily enhancing the mass fallback rate onto the SMBH by orders of magnitude as gas from the star is accreted (Hills, 1975; Rees, 1988). Non-thermal synchrotron emission from these tidal disruption events (TDEs) is best observed in the radio or millimeter and has been variously ascribed to jets, non-relativistic outflows, interactions between unbound stellar debris and the circumnuclear medium, or some combination thereof (see Alexander et al. 2020 for a review). A small fraction ($\lesssim 1$%) of known TDEs produce powerful relativistic jets with peak radio luminosities $\nu L_{\nu}\gtrsim 10^{41}$ erg s-1 (Zauderer et al., 2011; Cenko et al., 2012; Brown et al., 2015; Andreoni et al., 2022), but most TDEs lack such jets and are instead discovered as long-lasting transients in X-ray or optical surveys (e.g., Bade et al. 1996; Saxton et al. 2021; Sazonov et al. 2021; van Velzen et al. 2021; Hammerstein et al. 2023a; Yao et al. 2023). Targeted radio follow up of individual X-ray and optically discovered TDEs has revealed that some such TDEs exhibit lower-luminosity radio emission ($\nu L_{\nu}\lesssim 10^{39}$ erg s-1) within a few months of discovery (e.g., van Velzen et al. 2016; Alexander et al. 2016, 2017; Stein et al. 2021; Cendes et al. 2021a), but early follow up efforts typically ceased if no radio emission was detected within $6-12$ months, leaving the late-time radio properties of most TDEs unknown.
The first searches for TDE radio emission on longer timescales of months to years post-discovery were optimized to target initially-relativistic jets viewed off-axis, as the radio emission of such jets would be suppressed at early times due to relativistic beaming and only become visible after jet deceleration. These searches resulted in only non-detections, confirming that the rarity of powerful relativistic jets was a physical and not a selection effect; however, these studies were not sensitive enough to probe the existence of weaker outflows on similar timescales (Komossa, 2002; van Velzen et al., 2013; Bower et al., 2013). Recently, deeper wide-field radio transient searches have resulted in the first blindly selected radio TDE candidates (i.e., nuclear radio transients discovered independently, without prior detection of a counterpart at other wavelengths), suggesting that fainter radio emission in TDEs may be $\sim 10$ times more common than powerful jets (Anderson et al. 2020; see also Somalwar et al. 2023a; Dykaar et al. 2024). However, it is still not clear if these radio-selected nuclear transients probe the same underlying population as optically selected or X-ray selected TDEs.
This is the second of two papers comprising the first large-scale attempt to quantify the prevalence of fainter, late-rising radio emission in TDEs and understand the physical origin of this emission. The first of the two papers, Cendes et al. (2024a) (hereafter C24), presented the surprising result that late-rising radio emission is common in optically selected TDEs, with $\sim 40$% of the observed objects showing variable radio emission years post-discovery. This is comparable to — or possibly slightly higher than — the rate of TDEs that exhibit “prompt” radio emission (i.e., within 6 months of discovery), although the sample size of TDEs with early radio observations remains small and the occurrence rate of prompt radio emission in TDEs thus remains poorly constrained (Alexander et al., 2020). The delayed radio flares characterized by C24 and others (e.g., Horesh et al. 2021a, b; Cendes et al. 2022b; Hajela et al. 2024) are generally faint ($\nu L_{\nu}\lesssim 10^{39}$ erg s-1, with a few exceptions) and have been variously ascribed to non-relativistic outflows launched several hundred days post-disruption (Horesh et al., 2021a; Cendes et al., 2022b), to off-axis relativistic jets launched promptly (Matsumoto & Piran, 2023; Sfaradi et al., 2024), to initially-choked precessing relativistic jets that only escape after an initially misaligned accretion disk aligns with the black hole spin axis (Teboul & Metzger, 2023; Lu et al., 2024), or to outflowing material encountering changes in the density of the ambient circumnuclear medium (Matsumoto & Piran, 2024; Zhuang et al., 2024). While the physical origin of TDE radio emission remains unclear, the wide range of emission onset times — from a few days post-discovery to $>1000$ days post-discovery — may suggest that multiple mechanisms are responsible for producing low-luminosity radio emission in TDEs.
The multi-wavelength properties of TDEs may shed light on the origin of late-rising radio emission, as they give us independent methods to trace the evolution of the SMBH accretion rate with time. Mapping the accretion rate has been attempted in a variety of ways, most commonly by modeling TDEs’ optical/UV or X-ray light curves (e.g. Mockler et al. 2019; Saxton et al. 2021). If delayed radio brightenings are caused by newly-launched accretion-driven outflows, then we expect some correlation between the time when radio emission first appears and the instantaneous accretion rate onto the SMBH at that time. Conversely, if the delayed radio brightenings are powered by a promptly-launched outflow encountering changes in the ambient environment, we would not expect such correlations.
While the optical through X-ray emission from TDEs is generally broadly consistent with thermal emission from hot gas near the SMBH, the exact connection between the observed optical, UV, and X-ray light curves of TDEs and the accretion rate has been the subject of intense theoretical debate (Roth et al. 2020 and references therein). One open question is whether the mass fallback rate (i.e. the rate at which stellar debris approaches the black hole from the point of disruption) maps directly to the accretion rate (i.e. the rate at which material crosses the event horizon), or if the latter is significantly delayed relative to the former. While the mass fallback rate may be easier to constrain observationally in some cases, it is the accretion rate that is relevant for predicting the launch times of radio-producing outflows. For this work, we therefore consider two extremal possibilities: 1) that the stellar debris rapidly circularizes and forms an accretion flow, such that the bolometric optical-through-UV light curve directly traces the mass fallback rate (typically expected to exceed the Eddington limit before declining as $t^{-5/3}$; Rees 1988) and the instantaneous accretion rate onto the SMBH (accounting for a minimal viscous delay, e.g., Guillochon & Ramirez-Ruiz 2013; Mockler et al. 2019); or 2) that the peak accretion rate onto the SMBH is significantly delayed relative to peak optical light, either because the debris circularization is inefficient (in which case the optical peak may be powered instead by shocks from stream-stream collisions, e.g., Piran et al. 2015; Shiokawa et al. 2015; Lu & Bonnerot 2020) or because radiation pressure initially prevents even rapidly circularized debris from collapsing into a disk while the mass fallback rate is high, instead forming a hot, extended quasi-spherical pressure-supported envelope that powers the early optical emission while keeping the accretion rate across the event horizon low (Loeb & Ulmer, 1997; Guillochon et al., 2014; Metzger, 2022). In Case 1, X-rays produced near the SMBH are reprocessed into the optical band by more extended material, so the accretion rate is best traced in the optical/UV (Coughlin & Begelman, 2014; Guillochon et al., 2014; Metzger & Stone, 2016; Roth et al., 2016; Dai et al., 2018; Mockler et al., 2019). In Case 2, tracking the X-ray luminosity may provide a better method of estimating the instantaneous accretion rate, as thermal X-ray emission in these models originates in the innermost part of the accretion flow and thus indicates that matter is successfully approaching the SMBH regardless of the origin of the optical emission (Auchettl et al., 2017; Gezari et al., 2017; Metzger, 2022). In both cases, state transitions in the accretion flow are expected as the accretion rate changes, causing the production and quenching of outflows, but the timing of these transitions relative to peak optical light will be different. We compare the onset of any observed radio emission to the expected timing of state transitions in the accretion flow in both cases, focusing on the highly super-Eddington and highly sub-Eddington accretion phases that are most likely to power outflows (e.g., Giannios & Metzger 2011; De Colle et al. 2012).
In C24, we modeled the radio emission detected in optically selected TDEs at late times using an equipartition analysis, discussed its properties in the context of other radio observations of TDEs, and calculated the occurrence rate of late-time radio emission. Here, we discuss the multi-wavelength properties of radio-observed optical TDEs in more detail and attempt to discriminate among possible physical models for delayed radio emission. In Section 2 we introduce our sample and describe our new multi-wavelength observations and the archival data that we use for our subsequent analysis. We describe our light curve modeling and compare the radio properties of our sample to their multi-wavelength emission and derived properties in Section 3. We discuss the implications of our results in Section 4 and we conclude in Section 5.
2 Sample definition and data collection
This work was originally motivated by a search for radio emission associated with 91 known TDE candidates in the VLA Sky Survey (VLASS). This search resulted in three detections of transient radio emission associated with high-confidence TDEs and three detections of likely unrelated radio sources (Appendix A). Surprisingly, the detected TDEs included a dramatic rebrightening of the optical TDE ASASSN-15oi discovered in VLASS epoch 1 at $\sim 3$ years post-disruption (Horesh et al., 2021a; Hajela et al., 2024), motivating the deeper radio searches presented in C24 and this paper. A second optical TDE in our sample, AT2018hyz, appears as an even brighter radio transient in VLASS epoch 3 (Cendes et al. 2022b; C24).
To probe the physical mechanism or mechanisms behind the newly discovered late-rising radio emission in optically detected TDEs, we utilize multi-wavelength observations of a sample of TDEs discovered in optical surveys between 2014 and 2020 (Table 1). Our sample contains 31 TDEs at redshift $z\lesssim 0.1$ for which targeted radio follow up was obtained. As we wish to understand if late-time radio flares have a different physical origin than prompt radio emission in TDEs, our sample consists of the 24 TDEs observed by C24, most of which had no published radio detections prior to that work, plus an additional 7 TDEs from the literature previously reported to exhibit radio emission at early times. In addition to compiling optical, UV, and X-ray photometry of our TDEs from the literature, we also obtained new late-time X-ray observations of all TDEs in our sample. We briefly summarize all data used in this paper below and provide details of our new data reduction and analysis. A complete list of all references for previously-published data used in this work is given in Appendix B (Table 4), along with all new radio and X-ray photometry used for our analysis (Tables 5, 6). The complete multiwavelength dataset will also be made available as part of the Open mulTiwavelength Transient Event Repository (OTTER; Franz et al. in prep).
Table 1: Our Sample: Optically-selected TDEs at $z\lesssim 0.1$ with radio observations discovered between January 2014 and October 2020. TDEs in bold are those with newly discovered radio emission, presented in full in our companion paper, C24.
TDE Name Redshift Discovery Date of first $\Delta t$ of first radio Radio Optical
Date radio detection detection (rest-frame days) Type Type
ASASSN-14ae 0.0436 2014 Jan 25 2020 May 25 2216 Delayed TDE-H+He
DES14C1kia 0.162 2014 Nov 11 None None TDE-He
ASASSN-14li 0.02058 2014 Nov 22 2014 Dec 24 44 Prompt TDE-H+He
iPTF15af 0.07897 2015 Jan 15 None None TDE-H+He
ASASSN-15oi 0.0484 2015 Aug 14 2016 Feb 12 177 Delayed TDE-He
iPTF16axa 0.108 2016 May 29 None None TDE-H+He
PS16dtm (AT2016ezh) 0.0804 2016 Aug 12 2020 Jun 6 1287 Delayed TDE-H
iPTF16fnl (AT2016fnl) 0.016328 2016 Aug 29 2017 Jan 29 151 Prompt TDE-H+He
OGLE17aaj* 0.116 2017 Jan 2 2021 May 1 1417* Host* TDE-H+He
AT2017eqx 0.1089 2017 Jun 7 None None TDE-H+He
AT2018zr 0.071 2018 Mar 2 2022 Oct 16 1599 Delayed TDE-H+He
AT2018bsi* 0.051 2018 Apr 9 2022 Dec 8 1622* Host* TDE-H+He
AT2018dyb 0.018 2018 Jul 11 2021 May 4 1010 Delayed TDE-H+He
AT2018fyk 0.059 2018 Sep 8 2024 Feb 20 1880 Delayed TDE-H+He
AT2018hco 0.088 2018 Oct 4 2021 Jun 12 903 Delayed TDE-H
AT2018hyz 0.04573 2018 Nov 6 2021 Jun 11 934 Delayed TDE-H+He
AT2018lna 0.091 2018 Dec 28 None None TDE-H+He
ASASSN-19bt (AT2019ahk) 0.0262 2019 Jan 29 2019 Mar 3 39 Prompt TDE-H
AT2019azh 0.0222 2019 Feb 22 2019 Mar 15 21 Prompt TDE-H+He
AT2019dsg 0.0512 2019 Apr 9 2019 May 22 43 Prompt TDE-H+He
AT2019ehz 0.074 2019 Apr 29 2021 Jun 11 722 Delayed TDE-H
AT2019eve 0.064 2019 May 5 2021 Jun 11 723 Delayed TDE-H
AT2019qiz 0.0151 2019 Sep 19 2019 Sep 26 6 Prompt TDE-H+He
AT2019teq 0.0878 2019 Oct 20 2022 Oct 19 1008 Delayed TDE-H+He
AT2020pj* 0.068 2020 Jan 2 2022 Oct 28 964* Host* TDE-H+He
AT2020mot* 0.07 2020 Jun 14 2021 Jan 29 214* Host* TDE-H+He
AT2020neh 0.062 2020 Jun 19 2022 Nov 8 823 Delayed TDE-H+He
AT2020nov* 0.084 2020 Jun 27 2020 Oct 16 102* Host* TDE-H+He
AT2020opy 0.159 2020 Jul 8 2020 Oct 6 85 Prompt TDE-H+He
AT2020wey* 0.02738 2020 Oct 8 2021 Feb 13 125* Host* TDE-H+He
AT2020vwl 0.035 2020 Oct 10 2021 Feb 23 118 Prompt TDE-H+He
*indicates origin of radio emission is either ambiguous or unrelated to the TDE.
These TDEs showed only H lines in their initial classification spectra, leading to their labeling as TDE-H in some references. However, here we classify them as TDE-H+He because they also exhibit He emission at $\gtrsim 50-200$ days.
iPTF16fnl has been previously referred to in the literature as a “delayed” radio TDE due to several radio upper limits at early times (Horesh et al., 2021b). However, its first radio detection is $153$ $(151)$ days post-discovery in the observer (rest) frame, which places it in the “prompt” category as defined here (Section 2.1).
2.1 Radio: VLA, MeerKAT, ATCA, and literature data
In C24, we presented new radio observations at 2 – 6 years post-discovery of 24 optically-selected TDEs at $z\lesssim 0.1$ discovered between January 2014 and October 2020. These 24 TDEs are drawn from the samples defined by van Velzen et al. (2021) and Hammerstein et al. (2023a), plus a few additional events from the literature (Foley et al., 2015; Blanchard et al., 2017; Gromadzki et al., 2019; Frederick et al., 2020; Angus et al., 2022). Following C24, we also include previously-published radio observations and public archival data from the VLASS and VAST radio surveys (Lacy et al., 2020; Murphy et al., 2013) when available. For this work, we reanalyzed one observation of PS16dtm previously reported in C24, taken with NSF’s Karl G. Jansky Very Large Array (the VLA) on 2017 Aug 22 (372 days post-discovery) under program 16B-318 (PI: Alexander), and report a tentative new detection of $39\pm 10$ $\mu$Jy at 6 GHz. For AT2018fyk, we also include the recent radio detections reported by Cendes et al. (2024b).
In this paper, we add 7 TDEs from the literature with promptly-detected radio emission, to compare and contrast with the late-rising radio TDEs. Six of these (ASASSN-14li, ASASSN-15oi, AT2019azh, ASASSN-19bt, AT2019qiz, and AT2020opy) are included in the population analyses by van Velzen et al. (2021) and/or Hammerstein et al. (2023a); the remaining TDE is the youngest object in our sample and the first TDE classified as part of the second phase of the Zwicky Transient Facility survey (ZTF-II; Bellm et al. 2019a), AT2020vwl (Yao et al., 2023; Goodwin et al., 2023a). For AT2019qiz, we also include a subset of the VLA observations collected under programs 19A-013, 20A-372 (PI: Alexander), and 21A-303 (PI: Hajela), a public observation collected as part of the VLASS epoch 3, and a MeerKAT observation collected under program SCI-20220822-YC-01 (PI: Cendes). The AT2019qiz radio data used in our analysis are listed in Table 5 in Appendix B, and the complete radio dataset for AT2019qiz will be presented in future work (Franz et al. in prep).
All VLA data were reduced in CASA (McMullin et al., 2007) using standard procedures. For the MeerKAT observation, we used the standard calibrated MeerKAT pipeline images available via the SARAO Science Data Processor (SDP) [^1] https://skaafrica.atlassian.net/wiki/spaces/ESDKB/pages/338723406/. All flux densities were measured using the imtool package within pwkit (Williams et al., 2017), which fits a Gaussian within a small box centered on the pixel coordinates provided by the user. This Gaussian can either be fixed to be the same size and shape as the telescope beam for a point source fit, or allowed to vary freely for the case of extended emission. (By default imtool will compute the flux density both ways and return the model that provides a better fit/lower residuals, using a similar algorithm to the MIRIAD task imfit.) For virtually all of the sources in this paper, we do not expect significant contamination from extended host galaxy emission due to the relatively large distances and low star formation rates of the TDE hosts. We therefore use a point source fit, as indeed preferred for our images.
The full list of literature sources used to compile our radio dataset for each TDE in our sample is given in Appendix B (Table 4). For the purpose of our analysis, we classify each TDE’s radio emission as “Prompt:” first radio detection $<6$ months post-discovery (8 TDEs); “Delayed:” first radio detection $>6$ months post-discovery (12 TDEs), “None:” all radio observations to date resulted in non-detections (5 TDEs), or “Host:” persistent radio emission likely unrelated to the TDE (6 TDEs). The “Host” classification also includes TDEs where the origin of the observed radio emission is ambiguous (e.g., because the emission is too faint to definitively assess variability). We classify TDEs that exhibit multiple radio peaks (ASASSN-15oi and AT2019dsg) based on the onset time of the brightest radio emission episode. The radio classification of each TDE is given in Table 1.
2.2 X-rays: archival data and new Swift-XRT observations
All of the TDEs in our sample have received at least some X-ray follow-up, although the cadence and data quality vary significantly. We provide full details of the literature X-ray observations used in Table 4 in Appendix B. To systematically explore the late-time X-ray properties of our sample, we triggered new target-of-opportunity (TOO) observations with the X-ray Telescope (XRT, Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004) of all TDEs in Table 1 without existing late-time X-ray coverage, which cover the timescale $\sim 2-8$ yrs since discovery. We also re-analyzed all the available Swift-XRT data of TDEs in our sample not already reported in the literature, using the online automated tools released by the Swift-XRT team (Evans et al., 2009) [^2] https://www.swift.ac.uk/user_objects/ and custom scripts (Margutti et al., 2013). We first extracted a 0.3–10 keV count-rate light-curve with a minimum number of counts per bin allowing for dynamic binning. In many cases, these are TDEs with only faint X-rays or non-detections, so we assume the same spectral model for all TDEs for the flux calibration and fix $N_{H}=0$. We considered two spectral models: a constant non-thermal power-law spectrum with photon index $\Gamma=2$; and a constant thermal spectrum consisting of a black-body model with $k_{b}T=0.05\,\rm{keV}$. These spectral models and parameters are broadly representative of the hard (power-law) and soft (black-body) X-ray spectra of TDEs (Saxton et al., 2021; Guolo et al., 2023). For the power-law and black-body models, the average count-to-flux factors are $\approx 5\times 10^{-11}\rm{erg\,s^{-1}cm^{-2}}/\rm{c\,s^{-1}}$ and $\approx 3.3\times 10^{-11}\rm{erg\,s^{-1}cm^{-2}}/\rm{c\,s^{-1}}$, respectively. In the following we employ the power-law model flux calibration for all the TDE X-ray light-curves; we emphasize that the difference of a factor $<2$ between the two flux calibrations has no impact on any of our major conclusions. All newly-analyzed Swift XRT data are presented in Table 6 in Appendix B.
2.3 Optical: Archival spectroscopy and photometry
Optically selected TDEs typically show a common set of spectral features, including hydrogen and/or helium emission lines (Arcavi et al., 2014). All of the TDEs in our sample have optical spectral classifications previously reported in the literature (Foley et al., 2015; Blanchard et al., 2017; Gromadzki et al., 2019; Frederick et al., 2020; van Velzen et al., 2021; Angus et al., 2022; Hammerstein et al., 2023a; Yao et al., 2023). We divide our sample into three spectral types: TDEs with only hydrogen lines in their spectra (hereafter TDE-H), TDEs with both hydrogen and helium lines (TDE-H+He), and TDEs with only helium lines (TDE-He). However, the relative strengths of these lines vary both across TDEs (Arcavi et al., 2014; van Velzen et al., 2021) and as a function of time for specific TDEs (e.g. Nicholl et al. 2019). For this paper, we follow C24 and conservatively define a TDE H+He as a TDE that exhibited He lines at any phase of its evolution, in contrast to some recent literature. By this definition, our sample consists of 5 TDE-H, 24 TDE-H+He, and 2 TDE-He. We list the spectral classifications of our TDEs in Table 1. While most of these are consistent with the classifications reported in van Velzen et al. (2021) and Hammerstein et al. (2023a), three of the TDEs in our sample (ASASSN-14ae, AT2018zr, and AT2018hyz) initially showed only H lines in their spectra, but later developed He II emission at $\sim 50-200$ d post-discovery (Holoien et al., 2014; Short et al., 2020; Hung et al., 2019). We therefore initially classify them as TDE H+He for our analysis, but also consider the possibility that they may more naturally belong to the TDE-H group.
We also collected archival optical and ultraviolet photometry to perform more detailed light curve modeling of the TDEs in our sample, as described in Section 3. We exclude DES14C1kia from this analysis, as the optical dataset for this TDE remains unpublished. A complete listing of the data used and the original references for each dataset are given in Appendix B.
2.4 Host galaxy observations
Most of our TDEs also have well-studied host galaxies. For our analysis in Section 3.1, we use published galaxy stellar masses ($M_{\rm gal}$), generally calculated via broadband modeling of archival near-IR through UV photometry, and published stellar velocity dispersions ($\sigma*$) derived from high-resolution optical spectroscopy (Wevers et al., 2017, 2019a; Yao et al., 2023). We also use the optical $u-r$ colors of the TDE host galaxies in our sample as compiled by French et al. (2020) and references therein (van Velzen et al., 2021; Hammerstein et al., 2021; Yao et al., 2023).
3 Multi-Wavelength Analysis
In this section, we briefly consider basic host galaxy properties of the TDEs in our sample and we compare the optical, UV, and X-ray emission of our TDEs to their radio emission. We model the optical and UV light curves of our TDEs to calculate key parameters of each disruption (e.g., the SMBH mass and the mass of the disrupted star) and search for possible correlations with their radio properties. We then extract the temporal evolution of the mass accretion rate onto the SMBH, $\dot{M}$, and compare the observed radio and X-ray evolution to the timing of potential state transitions in the accretion flow based on our derived $\dot{M}$. Finally, we compare the radio and X-ray luminosities of our TDEs to those of AGN, to gain additional insight into $\dot{M}$ and search for possible similarities with the accretion states of those systems.
3.1 Optical spectroscopic classification, host properties, and SMBH masses
We first search for possible correlations between our radio classifications and the basic optical properties of the TDEs in our sample and their host galaxies. We find that 4 of the 12 TDEs that exhibit delayed radio emission have an optical spectral classification of TDE-H. For our sample of 31 objects containing 5 TDE-H total, the probability of at least 4 of the 5 TDEs also falling into this group by chance is $p=0.06$. Thus, based on the current dataset and our conservative definition of a TDE-H as one without any He emission at any epoch, we find no statistically significant correlation between the timing of radio emission and optical spectral type. However, interestingly, if the three TDEs for which He lines only emerge at $t>50$ d are reclassified as TDE-H rather than TDE-H+He, then our sample has a total of 8 TDE-H, of which 7 exhibit delayed radio emission. In this case, the probability of chance coincidence decreases to $p=0.002$.
Previous studies of the optical $u-r$ colors of TDE host galaxies have shown that TDE hosts are preferentially located in the “green valley” (French et al. 2020 and references therein; Hammerstein et al. 2021; Yao et al. 2023). We find no obvious correlation between the bulk host galaxy classifications (quiescent vs. star forming; red vs. green vs. blue) and the type of radio emission in our TDE sample (prompt vs. delayed vs. none/host). We defer further investigation of possible impacts of the host environment on the radio properties of TDEs to future work.
Finally, we compute the SMBH masses for all TDEs in our sample. As SMBH masses previously reported in the literature are derived using several different methods, here we use two independent techniques: a transient-independent method based on observations of the TDE host galaxies ($M_{\rm BH,host}$) and a TDE-only method based on fitting the TDE optical and UV light curves ($M_{\rm BH,mosfit}$). First, we use archival host galaxy observations to self-consistently recalculate the SMBH masses for all of our TDEs and present them in Table 2, along with the original measurement used to derive each value. For galaxies with published stellar velocity dispersions, we calculate the black hole mass using the $M_{\rm BH}-\sigma*$ relation from Kormendy & Ho (2013). For galaxies that do not have published velocity dispersions but do have published stellar masses, we use the $M_{\rm BH}-M_{\rm gal}$ relation for TDE host galaxies derived by Yao et al. (2023), which is calibrated against the Kormendy & Ho (2013) relation. [^3] OGLE17aaj’s host galaxy does not have a published $\sigma*$ measurement or $M_{\rm gal}$ value; we therefore use the SMBH mass estimate from Gromadzki et al. (2019), $\log(M_{\rm BH}/M_{\odot})=7.37$.
TDE light curves have also been used to estimate SMBH masses directly (Mockler et al., 2019; Ryu et al., 2020; Sarin & Metzger, 2024; Mummery et al., 2024b). The masses derived in this way are typically found to be consistent with SMBH masses derived from the host galaxy properties, with some scatter. We choose to model the optical and UV light curves of our TDEs using the Modular Open Source Fitter for Transients (MOSFiT; Guillochon et al. 2018; Mockler et al. 2019) to obtain a second, independent estimate of the SMBH mass for each TDE. For our sample, we find that the two methods yield generally consistent results, although on average the MOSFiT-derived masses are slightly higher (Figure 1, left). This is similar to the results obtained by Mockler et al. (2019) for their TDE sample. For our analysis, we do not make any claims as to which method of deriving the SMBH mass is more accurate; instead, we consider both sets of masses independently whenever possible. We discuss our MOSFiT modeling further in the next two sections.
Table 2: SMBH masses for the TDEs in our sample derived from properties of their host galaxies. The $M_{\rm BH,host}$ values have been computed using the $M_{\rm BH}-\sigma*$ relation of Kormendy & Ho (2013) when possible, or using the $M_{\rm gal}-M_{\rm BH}$ relation from Yao et al. (2023) otherwise. The source used for each galaxy stellar mass or $\sigma*$ measurement is indicated in the Reference column.
TDE Name $\log(M_{\rm gal}/M_{\odot})$ $\sigma*$ (km s-1) Reference $\log(M_{\rm BH,host}/M_{\odot})$
ASASSN-14ae $53\pm 2$ Wevers et al. (2019b) $6.0\pm 0.3$
DES14C1kia $10.1$ Wevers et al. (2019b) $6.6$
ASASSN-14li $81\pm 2$ Wevers et al. (2019b) $6.8\pm 0.3$
iPTF15af $106\pm 2$ Wevers et al. (2019b) $7.3\pm 0.3$
ASASSN-15oi $61\pm 7$ Wevers et al. (2019b) $6.2\pm 0.4$
iPTF16axa $82\pm 3$ Wevers et al. (2019b) $6.8\pm 0.3$
PS16dtm $45\pm 3$ Xiao et al. (2011) $5.7\pm 0.4$
iPTF16fnl $55\pm 2$ Wevers et al. (2019b) $6.0\pm 0.3$
OGLE17aaj* Gromadzki et al. (2019) $7.37$
AT2017eqx $9.36^{+0.1}_{-0.1}$ Nicholl et al. (2019) $5.3\pm 0.4$
AT2018zr $49.79\pm 4.93$ Hammerstein et al. (2023b) $5.8\pm 0.4$
AT2018bsi $117.54\pm 8.12$ Hammerstein et al. (2023b) $7.5\pm 0.3$
AT2018dyb $96\pm 1$ Leloudas et al. (2019) $7.1\pm 0.3$
AT2018fyk $158\pm 1$ Wevers (2020) $8.0\pm 0.3$
AT2018hco $10.03^{+0.12}_{-0.16}$ Hammerstein et al. (2023a) $6.5\pm 0.3$
AT2018hyz $66.62\pm 3.12$ Hammerstein et al. (2023b) $6.4\pm 0.3$
AT2018lna $36.43\pm 4.52$ Hammerstein et al. (2023b) $5.2\pm 0.4$
ASASSN-19bt $10.04^{+0.34}_{-0.04}$ Holoien et al. (2019a) $6.5\pm 0.4$
AT2019azh $68.01\pm 2.02$ Hammerstein et al. (2023b) $6.4\pm 0.3$
AT2019dsg $86.89\pm 3.92$ Yao et al. (2023) $6.9\pm 0.3$
AT2019ehz $46.65\pm 11.83$ Hammerstein et al. (2023b) $5.7\pm 0.6$
AT2019eve $9.26^{+0.11}_{-0.17}$ van Velzen et al. (2021) $5.1\pm 0.4$
AT2019qiz $71.85\pm 1.93$ Hammerstein et al. (2023b) $6.5\pm 0.3$
AT2019teq $9.95^{+0.07}_{-0.11}$ Hammerstein et al. (2023a) $6.3\pm 0.3$
AT2020pj $10.01^{+0.07}_{-0.08}$ Yao et al. (2023) $6.4\pm 0.3$
AT2020mot $76.61\pm 5.33$ Yao et al. (2023) $6.7\pm 0.3$
AT2020neh $40\pm 6$ Angus et al. (2022) $5.4\pm 0.5$
AT2020nov $127\pm 28$ Earl et al. (2024) $7.6\pm 0.5$
AT2020opy $10.01^{+0.13}_{-0.14}$ Hammerstein et al. (2023a) $6.4\pm 0.3$
AT2020wey $53.54\pm 4.75$ Hammerstein et al. (2023b) $6.0\pm 0.4$
AT2020vwl $48.49\pm 2.00$ Yao et al. (2023) $5.8\pm 0.4$
*OGLE17aaj’s host galaxy does not have a published stellar mass or velocity dispersion; we therefore assume the value of $\log\left(M_{\rm BH}/M_{\odot}\right)\sim 7.37$ computed by Gromadzki et al. (2019).
3.2 MOSFiT modeling: physical parameters
We fit the optical and UV photometry for the TDEs in our sample with MOSFiT (Guillochon et al., 2018), using version 1.1.9 of the TDE module first developed by Mockler et al. (2019) following Guillochon & Ramirez-Ruiz (2013); Guillochon et al. (2014). The original references for the included data are given in Table 4 in Appendix B. We use the same setup, priors, and model assumptions as those described in Nicholl et al. (2022). In particular, MOSFiT assumes that mass falling onto the SMBH accretes promptly, such that the observed optical luminosity is directly proportional to the mass fallback rate and the accretion rate, $L=\epsilon\dot{M}c^{2}$ (where $\epsilon$ is the radiative efficiency). [^4] A possible viscous delay is included in the model, but when applied to our observations this delay is generally found to be very short (consistent with zero in most cases), in line with previous work (Mockler et al., 2019; Mockler & Ramirez-Ruiz, 2021; Nicholl et al., 2022). This corresponds to Case 1 as defined in the introduction. This assumption likely breaks down at very late times, when TDEs’ optical/UV emission often exhibits a plateau phase that is not well-captured by MOSFiT but can be better modeled by a bare accretion disk (van Velzen et al., 2019a; Mummery et al., 2024a, b). For the TDEs in our sample that show optical/UV plateaus, we therefore only include photometry prior to the onset of the plateau phase in our modeling, as detailed in Appendix B.
We are able to find a converged model solution for all TDEs in our sample except for DES141Ckia (which has no published photometry) and AT2018fyk (whose double-peaked light curve is not well-modeled by MOSFiT’s one-zone model assumptions). A complete description of the free parameters of our model, together with the best-fit values and $1\sigma$ confidence intervals for the physically relevant parameters are given in Appendix C. We first use these results to search for possible correlations between the radio properties of our TDEs and other physical parameters of the disruption. We divide our sample into three groups for this analysis: those with radio type “Prompt” in Table 1 (i.e., TDEs that show early/prompt radio emission), “Delayed” TDEs in Table 1 (those that show late/delayed radio emission), and “None” or “Host” TDEs in Table 1 (those that show no radio emission or only host or faint ambiguous radio emission). We also considered the None and Host TDEs as two independent categories, but this does not significantly change our results. We then performed an Anderson-Darling (AD) test [^5] The Anderson-Darling test is similar to the Kolmogorov–Smirnov (KS) test used by some previous authors (e.g., Nicholl et al. 2022), but it has been shown to have added sensitivity to the tails of the distributions in datasets with large numbers of free parameters (Hou et al., 2009). We note that a KS test applied to our data gives similar results. for each of the parameters listed in Appendix C (Table 7) plus $M_{\rm BH,host}$ (the SMBH masses derived from host galaxy properties presented in Table 2).
In almost all cases, we find no obvious differences in the parameter distributions between any two of the three radio classes (see Table 8 in Appendix C for the full results). The one exception is the photospheric radius normalization parameter $R_{\rm ph,0}$, which is larger for the delayed radio TDEs than for the prompt radio TDEs ($p=0.026$). There is also a marginal suggestion ($p\sim 0.05$) that the TDEs with delayed radio emission have lower SMBH masses than either the TDEs with prompt radio emission or than the host+none group – but only when $M_{\rm BH,host}$ is used; no significant differences are found between the $M_{\rm BH,mosfit}$ distributions for these groups. We plot the distributions of $R_{\rm ph,0}$, $M_{\rm BH,host}$, and $M_{\rm BH,mosfit}$ in Figure 1. While achieving three $p$ values $\lesssim 0.05$ out of 36 statistical tests performed is not a highly significant result, in Section 4 we briefly explore the possible implications if these correlations are verified with a larger TDE sample in future work. We emphasize that null results in our current sample do not necessarily imply that correlations between the radio properties and the various parameters tested are not present, given the limitations of existing data. For example, several of the TDEs that we classify as radio type Delayed have no radio observations at $t<6$ months, meaning that we have no way of knowing if they also displayed prompt radio emission and are thus misclassified here, potentially diluting any underlying trends (Section 4.2.3).
Figure 1: Left: The host-derived BH masses $M_{\rm BH,host}$ compared to the MOSFiT-derived BH masses $M_{\rm BH,mosfit}$ for our sample. The dashed line shows the 1:1 relation. The marker color indicates the timing of radio emission for each TDE and the symbol indicates its optical spectral type. Right: $M_{\rm BH,host}$ compared to the MOSFiT-derived normalized photospheric radii $R_{\rm ph,0}$. Delayed radio TDEs (yellow points) exhibit larger $R_{\rm ph,0}$ than Prompt radio TDEs (dark blue points) and are more likely to have optical spectral type TDE-H (stars).
Figure 2: $\dot{M}$ curves from MOSFiT (solid lines; shaded region shows $1\sigma$ uncertainties) together with the radio (dark blue points) and X-ray (light blue points) light curves of TDEs where the approximate launch time of the outflow can be calculated from analysis of the full multi-frequency radio dataset. We normalize $\dot{M}$ by $\dot{M}_{\rm Edd}$ using both $M_{\rm BH,host}$ (orange) and $M_{\rm BH,mosfit}$ (black). Horizontal lines show the values of $\dot{M}$ corresponding to possible state transitions in the accretion flow; accretion-driven outflows are most likely for $\dot{M}\gtrsim 0.3\dot{M}_{\rm Edd}$ or $\dot{M}\lesssim 0.03\dot{M}_{\rm Edd}$. The vertical red lines (shaded bars) show the estimated launch times (and associated uncertainties) of the late-rising radio outflows as described in the text.
Figure 3: Same as Figure 2 for the remaining TDEs in our sample with a delayed radio emission episode and converged MOSFiT models. For six of these TDEs, C24 provide a rough estimate of the outflow launch time (dashed vertical red lines). For AT2019dsg, the dashed line indicates the possible launch time of the second radio flare. For PS16dtm, the open circle denotes a marginal possible detection discovered in our reanalysis of the data (Section 2.1). C24 do not estimate a launch time for AT2020neh due to its faintness.
Figure 4: Same as Figures 2 and 3 for our remaining TDEs previously found to exhibit prompt radio emission (within 6 months of discovery). Unlike the TDEs in Figures 2 and 3, several of these TDEs require outflow launch dates corresponding to the initial rise of the optical light curve rather than to a period of super-Eddington accretion (e.g., Goodwin et al. 2023a; Christy et al. 2024). This may suggest a different mechanism is required to power early radio outflows in TDEs.
3.3 MOSFiT modeling: accretion rate
An additional derived output of the MOSFiT code is a plot of how the accretion rate onto the SMBH, $\dot{M}$, varies with time. To further explore how the radio behavior of TDEs may correlate with the SMBH accretion rate, we use MOSFiT to extract the temporal evolution of $\dot{M}$ in our TDEs and compare their observed radio and X-ray evolution to the timing of potential state transitions in the accretion flow (i.e., the times when the accretion disk may switch in or out of a state capable of powering outflows). We assume here that outflows are most likely for super-Eddington accretion ($\dot{M}\gtrsim 0.3\;\dot{M}_{\rm Edd}$ and highly sub-Eddington accretion $\dot{M}\lesssim 0.03\;\dot{M}_{\rm Edd}$).
In Figures 2-4 we show the $\dot{M}$ curves together with the radio and X-ray light curves for all TDEs from our sample with unambiguously transient radio emission and converged MOSFiT models666i.e., all TDEs with radio type “Prompt” or “Delayed” in Table 1 except for AT2018fyk, whose optical emission cannot be well-fit by MOSFiT. Further late-time radio monitoring of the TDEs with radio types Host or None is ongoing and will be presented in future work.. Specifically, Figure 2 shows TDEs for which we have a reasonably well-constrained estimate of the outflow launch time [^7] ASASSN-14ae, AT2018hco, AT2018hyz, and AT2019eve have multiple radio SEDs at late times, so it is possible to use the evolution of the derived physical size of the radio-emitting region to obtain a rough estimate of the time at which their delayed outflows were launched (Cendes et al. 2022b and C24). Similarly, Horesh et al. (2021b) model the multi-frequency radio light curves of iPTF16fnl to constrain the likely radio outflow launch time to $\sim 56-153$ days post-discovery. Multi-wavelength observations of ASASSN-15oi suggest that an outflow launched at the time of the soft X-ray peak may provide a good explanation for its second radio flare (Horesh et al., 2021a; Hajela et al., 2024)., Figure 3 shows TDEs with delayed radio emission where we have at most an approximate constraint on the outflow launch time from C24888plus AT2019dsg, for which we show the approximate onset time of its secondary delayed peak from C24, and AT2020neh, which is too faint for an accurate launch estimate., and Figure 4 shows the remaining TDEs with prompt radio emission. We normalize $\dot{M}$ to $\dot{M}_{\rm Edd}$, where $\dot{M}_{\rm Edd}$ is the accretion rate required to power emission equal to the Eddington luminosity of the SMBH assuming a radiative efficiency $\epsilon$; $L_{\rm Edd}=\epsilon\dot{M}_{\rm Edd}c^{2}$. We use the radiative efficiency from MOSFiT for each TDE and our two independent estimates of the SMBH mass from Section 3.1. We find that most TDEs in our sample achieve a peak $\dot{M}$ that approaches or mildly exceeds $\dot{M}_{\rm Edd}$. The peak radio luminosity of a TDE does not appear to correlate with the maximum $\dot{M}/\dot{M}_{\rm Edd}$ achieved onto its SMBH.
We next investigate if the onset time of the radio emission in our TDEs correlates with any specific $\dot{M}/\dot{M_{\rm Edd}}$ ratio, beginning with the TDEs in Figure 2 where we have the best constraints. We find that three of these TDEs (AT2018hco, ASASSN-15oi, and iPTF16fnl) have outflows that were plausibly launched when the accretion rate was a few percent of $\dot{M}_{\rm Edd}$, although for the latter two objects we are limited by the systematic uncertainties on the SMBH masses. ASASSN-15oi’s and iPTF16fnl’s radio emission may also have been launched earlier, possibly when $\dot{M}$ was still super-Eddington. [^9] This is particularly likely for ASASSN-15oi, whose multi-wavelength observations suggest that accretion in this TDE may be delayed a few hundred days relative to our MOSFiT prediction (Hajela et al., 2024). AT2018hyz and AT2019eve’s radio outflows do not appear until several hundred days after the $0.03\dot{M}/\dot{M}_{\rm Edd}$ transition, implying that either accretion is delayed relative to the MOSFiT prediction for these two events, or that some other process is responsible for their radio brightenings. [^10] The launch time from Cendes et al. (2022b) assumes that AT2018hyz’s radio outflow is non-relativistic; if its radio emission is instead due to a relativistic off-axis jet, then a much earlier launch time coinciding with the predicted period of super-Eddington accretion is possible (Matsumoto & Piran, 2023; Sfaradi et al., 2024). Finally, ASASSN-14ae’s very late outflow launch time is difficult to reconcile with either state transition, assuming the outflow launch time and the mass of its SMBH are both correct. This may suggest that either accretion disk formation is significantly delayed in this TDE relative to the prediction from MOSFiT, or some other unrelated process is required to explain its radio rise.
Although the onset time of radio emission is less clear for the TDEs in Figure 3, we can roughly approximate the launch time of each outflow based on the radio light curve (C24; dashed red lines in Figure 3). [^11] We note that our new tentative detection of PS16dtm (Section 2.1; open circle in Figure 3) pre-dates the launch time suggested by C24 for this TDE; it is therefore likely that its radio outflow was launched significantly earlier ($\sim 1$ year post-discovery). The systematic uncertainties on the SMBH masses and, in some cases, the sparse temporal sampling of the radio light curves preclude a definitive association of the radio emission with a particular accretion rate for these TDEs. It is nevertheless allowed within the uncertainties that the delayed radio emission in all of these events (including the secondary peak in the light curve of AT2019dsg) is driven by a state change in the accretion flow associated with $\dot{M}$ declining to $\lesssim 0.03\dot{M}_{\rm Edd}$. Again, in most cases, peak accretion onto the SMBH would have to be delayed by several hundred days or more relative to the MOSFiT predictions to associate the outflow launch times with super-Eddington accretion instead. [^12] PS16dtm and AT2019ehz are the two possible exceptions, if their host galaxy-derived SMBH masses rather than their MOSFiT-derived SMBH masses are assumed to be correct.
Finally, Figure 4 shows the $\dot{M}$ evolution and radio light curves of the remaining 6 TDEs in our sample for which prompt radio emission was detected within 6 months of discovery. For most of the Prompt radio TDEs, the first epoch of radio follow up unfortunately does not occur until the time of peak optical light or later, as there is typically a delay between transient discovery and the timescale on which most optical TDEs are classified and publicly announced to the community. While direct evidence for pre-peak radio emission in these TDEs is therefore limited, AT2019qiz’s first radio detection (7 days post-discovery, the earliest radio detection of any TDE in our sample) occurs when $\dot{M}$ is sub-Eddington and has not yet peaked, suggesting that the process powering prompt radio emission in this TDE may not be related to a high accretion rate onto the SMBH. Additionally, for several events (ASASSN-19bt, AT2019dsg, AT2020vwl), equipartition analyses of their multi-frequency radio SEDs suggest that the launch times of their radio outflows coincide with the time their optical light curves began to rise – well before $\dot{M}$ peaks (Cendes et al., 2021a; Goodwin et al., 2023a; Christy et al., 2024).
3.4 X-ray light curves and the Fundamental Plane
Figure 5: The quasi-simultaneous ($\delta t/t<0.25$) X-ray and radio luminosities of our TDEs shown relative to the Fundamental Plane relationship (dashed black line with intrinsic scatter given by the shaded gray region) and AGN sample (gray X’s) from Merloni et al. (2003). Arrows indicate upper limits and points from the same TDE are connected by a solid line. The color of each point indicates the time in days since discovery (dark blue = earliest, yellow = latest). We highlight six particularly noteworthy TDEs with unique symbols given in the figure legend. We find that X-ray and radio-detected TDEs generally follow the fundamental plane at early times (within 6 months of discovery), as do a subset of the TDEs with radio detections at late times. ASASSN-15oi (stars) is consistent with the plane during its first radio flare but deviates from the plane during its second radio flare, possibly suggesting different physical origins for the radio emission on these two timescales.
The X-ray properties of our TDEs may also connect to the instantaneous accretion rate onto the SMBH (e.g., Metzger 2022). Together with our SMBH mass measurements (Section 3.1), we can use our radio and X-ray light curves to track the evolution of individual TDEs in the $L_{X}$-$L_{R}$ phase space over thousands of days and compare the properties of our sample to the Fundamental Plane of black hole activity. The Fundamental Plane is a well-established correlation between radio luminosity, X-ray luminosity, and black hole mass that holds across many orders of magnitude for black holes accreting at low rates (Merloni et al., 2003; Falcke et al., 2004). If delayed radio emission is produced in TDEs after the accretion rate has dropped to similarly low levels, we may expect these TDEs to obey the same relation at late times.
We show our TDE sample on the Fundamental Plane in Figure 5, using the SMBH masses derived from the host galaxy properties [^13] Using $M_{\rm BH,host}$ rather than $M_{\rm BH,mosfit}$ allows us to include DES14C1kia and AT2018fyk, which do not have converged MOSFiT models. (Table 2). The original sample of AGN from Merloni et al. (2003) are also shown for comparison. As this relation is calibrated using radio observations collected at 5 GHz, we restrict ourselves to radio data taken at similar frequencies ($\nu<10$ GHz). We furthermore only show TDEs for which quasi-simultaneous radio and X-ray data are available, i.e., observations for which $\delta t/t<0.25$. For TDEs with high-cadence observations, if multiple radio/X-ray data point pairs fall within this tolerance, we select the data pairs with the smallest separation in time and frequency space, such that each available data point is used at most once in the figure. We highlight six TDEs of particular interest in the figure legend. Among the TDEs with sufficiently good radio and X-ray coverage, we see that all of our prompt radio TDEs and many of our delayed radio TDEs are consistent with the Fundamental Plane. Conversely some TDEs, such as ASASSN-15oi, move away from the Plane at late times.
4 Discussion
We now attempt to synthesize our multi-wavelength analyses into a physical model, or models, to explain delayed radio emission in TDEs. Any model must be able to explain our key observational findings from Section 3, which we briefly summarize here:
  • TDEs with delayed radio emission are more likely to entirely lack He emission in their optical spectra or to only exhibit it at late times, in comparison to TDEs with prompt radio emission ($p=0.002$).
  • TDEs with delayed radio emission have larger photospheric radii than TDEs with prompt radio emission ($p=0.026$).
  • Some TDEs (e.g., AT2018hco, the second radio flare in AT2019dsg) have radio outflow launch times that roughly coincide with MOSFiT’s predicted transition to a highly sub-Eddington state $\dot{M}<0.03\dot{M}_{Edd}$. However, we find that other delayed radio outflows are launched up to several hundred days later than the MOSFiT predictions for the timing of this state transition (e.g., in ASASSN-14ae).
  • Similarly, while the radio launch times of a few TDEs in our sample may be marginally consistent with MOSFiT’s predicted super-Eddington phase $\dot{M}>0.3\dot{M}_{Edd}$ (ASASSN-15oi’s first flare, iPTF16fnl, PS16dtm, AT2019ezh), the radio emission in most of our delayed radio TDEs can only be explained as outflows powered by super-Eddington accretion if the super-Eddington phase is delayed by several hundred days relative to the MOSFiT prediction.
  • The onset of prompt radio emission in some TDEs precedes the period of peak accretion onto the SMBH.
  • Many TDEs, including all TDEs in our sample with prompt radio emission and some TDEs with delayed radio emission, are broadly consistent with the Fundamental Plane of black hole activity.
  • The two TDEs that show the most dramatic late-time radio brightenings (AT2018hyz and ASASSN-15oi’s second flare) are both inconsistent with the Fundamental Plane at late times.
Broadly speaking, there are two classes of models that have been proposed to explain radio emission in TDEs: models in which the radio emission is directly powered by accretion-driven outflows (e.g. Strubbe & Quataert 2009; Giannios & Metzger 2011; De Colle et al. 2012; Alexander et al. 2016; van Velzen et al. 2016; Metzger 2022), and those where it originates from some other aspect of the disruption process unrelated to SMBH accretion. The latter category includes emission from shocks between the unbound portion of the stellar debris stream and the ambient environment (Krolik et al., 2016), emission from outflows launched by stream-stream collisions during the circularization of the bound stellar debris (Lu & Bonnerot, 2020), and sudden brightenings of a shock that encounters density variations in the circumnuclear environment (Zhuang et al., 2024). In the first category, we might expect correlations between the radio behavior and other probes of the accretion rate (such as the optical or X-ray emission), while in the second category we would not expect such correlations. As we explore in this section, the existence of the various multi-wavelength connections uncovered in our analysis thus suggest that the late-time radio properties of our sample are driven by accretion onto the SMBH, rather than by environmental interactions.
4.1 Overview of accretion-driven jets and outflows in TDEs
As discussed in Giannios & Metzger (2011), there are two phases after a TDE where an accretion-driven relativistic jet or other radio-producing outflow may be formed. Firstly, if the process of debris circularization and disk formation is sufficiently rapid, then a jet may accompany the early super-Eddington phase ($\dot{M}\gg\dot{M}_{\rm Edd})$, in analogy with gamma-ray bursts. Such an early, super-Eddington jetted phase is likely responsible for highly energetic jetted TDEs such as Swift J1644+57 (De Colle et al., 2012). This early jet phase ends once the disk becomes geometrically thin at $\dot{M}\lesssim 0.3\dot{M}_{\rm Edd}$ and can no longer retain its magnetic flux (e.g., McClintock et al. 2006; Tchekhovskoy et al. 2014). This transition typically occurs hundreds of days post-disruption, consistent with the observed shut-off time of the X-rays in jetted TDEs (Zauderer et al., 2013; Cenko et al., 2012; Eftekhari et al., 2024). Super-Eddington accretion has also been proposed to power non-relativistic winds in TDEs that produce low-luminosity radio emission at early times (e.g., Strubbe & Quataert 2009; Alexander et al. 2016). The presence of a wind may explain why only powerful super-Eddington jets have been seen at early times: if the star’s orbital plane is initially misaligned with the black hole spin axis, the jet will precess and less powerful jets may only be able to break out of the wind after the disk undergoes sufficient Lense-Thirring precession to come into alignment with the spin (Teboul & Metzger, 2023; Lu et al., 2024). This will then delay any radio emission from less powerful super-Eddington jets. Radio emission from powerful relativistic jets may also appear delayed if the jets are viewed off-axis, as relativistic beaming will suppress the emission at early times.
Alternatively, a jet may form at very late times, once the accretion rate decreases to $\dot{M}\lesssim 0.03\dot{M}_{\rm Edd}$ (Giannios & Metzger, 2011). When the accretion rate becomes this low, the accretion flow is predicted to transition back to radiatively inefficient and geometrically thick, which can again power outflows (e.g., Narayan & Yi 1995; Blandford & Begelman 1999). Such behavior is supported by detections of radio jets from X-ray binaries (XRBs) in the low-hard state and bright radio emission consistent with compact jets from AGN at low Eddington ratios (e.g., Fender et al. 1999; Falcke et al. 2004; Sikora et al. 2007; Corbel et al. 2013). The rapid X-ray and optical spectral and photometric variability observed in changing-look AGN (Matt et al., 2003; Ricci & Trakhtenbrot, 2023) also supports the existence of this state change, as this variability is consistent with the sudden appearance or disappearance of a jet and changing-look AGN are predominantly observed to have Eddington ratios of a few percent, right at this expected transition. (This is much lower than the typical Eddington ratios seen in the general AGN population; MacLeod et al. 2019; Green et al. 2022; Guo et al. 2024.) A delayed radio jet was recently discovered in a changing-look AGN $\sim 6$ years after its original optical outburst, providing direct evidence that these events can power outflows (Meyer et al., 2025). In TDEs, this late jet phase is expected to occur on timescales of years to decades.
We expect our TDEs to adhere to the Fundamental Plane of black hole activity (Section 3.4) during periods where their SMBHs are accreting at a low fraction of the Eddington rate (Falcke et al., 2004; Jonker et al., 2020). Conversely, if the accretion rate is still high during some observed delayed radio brightenings, then we might expect the radio and X-ray emission to deviate from the Fundamental Plane for these TDEs. We may also expect to see a given TDE evolve along the Plane or towards/away from it as $\dot{M}$ changes and the accretion flow possibly undergoes state changes. The radio and X-ray emission of some TDEs also may not follow the Fundamental Plane at any epoch if the radio emission is not directly driven by accretion onto the SMBH (e.g., if prompt emission is produced instead by the shock between the unbound portion of the stellar debris stream and the ISM; Krolik et al. 2016, and/or if late-time emission is due to a previously-launched outflow encountering a change in the ambient environment; Zhuang et al. 2024). In the case of radio emission dominated by a precessing jet whose escape is delayed due to an initial misalignment between the disk and the SMBH spin, we also do not expect any clear correlation between the simultaneous radio and X-ray emission, as the radio onset time is determined by the alignment timescale rather than the accretion rate in these models (Teboul & Metzger, 2023; Lu et al., 2024).
4.2 Origin(s) of the delayed radio emission
Our delayed radio TDEs exhibit radio emission on a wide range of timescales (hundreds to thousands of days post-discovery) and with a wide range of luminosities, suggesting that a single mechanism may not be sufficient to explain the entire sample. In this section, we discuss the various types of accretion-driven outflows that may contribute to the delayed radio emission. As described in Section 3, we constrain the accretion rate as a function of time for the TDEs in our sample via two methods: modeling the optical/UV light curves using MOSFiT, and comparing their X-ray and radio properties directly to the Fundamental Plane of black hole activity.
4.2.1 (Delayed) super-Eddington accretion
We first consider scenarios in which the delayed radio emission observed in our TDE sample is powered by non-relativistic outflows launched during a phase of super-Eddington accretion. While this is unlikely to explain TDEs with radio emission delayed by thousands of days (e.g. ASASSN-14ae) [^14] It is possible to delay peak accretion onto the SMBH by this much under certain assumptions (for example in the models for low-mass SMBHs explored by Guillochon & Ramirez-Ruiz 2015, where the debris stream may only self-intersect after dozens of pericenter passes and debris circularization can therefore be delayed by years). However, delaying the stream-stream collisions and debris circularization phases to this extent also spreads out the resulting mass fallback such that the peak accretion rate is likely no longer super-Eddington, and the optical evolution is likely much slower than observed in ASASSN-14ae., it may be a plausible explanation for TDEs where a radio outflow is launched hundreds days after discovery.
Many of our delayed radio TDEs have larger optical photospheric radii at early times and delayed or absent He lines. Previous work by Nicholl et al. (2022) found that TDE-H may arise from either less complete stellar disruptions (i.e., those with a smaller impact parameter, $b$) or from those in which an accretion disk does not form promptly. Our observed correlation between delayed radio emission and a lack of He emission could thus suggest that less complete stellar disruptions are also more likely to produce delayed radio emission, but we disfavor this interpretation as we find no statistically-significant difference in the $b$ values of our Prompt and Delayed TDE sub-samples (Appendix C). Instead, the fact that our delayed radio subsample both lacks prompt He emission and has larger optical photospheres favors the delayed accretion hypothesis. Nicholl et al. (2022) also found that their TDE-H exhibit larger photospheric radii than other types of TDEs, consistent with previous studies (Hinkle et al., 2020; van Velzen et al., 2021; Charalampopoulos et al., 2022). Nicholl et al. (2022) suggest that this could imply the early optical emission in TDE-H could be powered by collision-induced outflows rather than promptly formed accretion disks, such that accretion onto the SMHB is delayed relative to peak optical light. This would also naturally delay the onset of radio emission produced by accretion-powered outflows.
If the optical and UV emission around peak light is produced by energy dissipated at the debris stream self-intersection point, as suggested by e.g., Piran et al. (2015); Jiang et al. (2016), then larger photospheric radii would correspond to TDEs where the self-intersection point is located further away from the SMBH. Nicholl et al. (2022) suggest that this could occur in TDEs around less massive SMBHs. In such TDEs, it seems plausible that it would take correspondingly longer for the debris to reach the event horizon, resulting in delayed peak accretion onto the SMBH. Our weak correlation between delayed radio emission and larger $R_{\rm ph,0}$ values (and our even more tentative correlation between delayed radio emission and smaller $M_{\rm BH,host}$ values) could then suggest that at least some delayed radio outflows in TDEs are accretion-driven, and specifically are powered by a phase of high (super-Eddington) accretion delayed several hundred days relative to peak optical light. [^15] While our sample also exhibits the inverse correlation between $R_{\rm ph,0}$ and SMBH mass seen by Nicholl et al. (2022) (Figure 1, right), it is interesting to note that this correlation is not seen in the recent flux-limited TDE sample studied by Yao et al. (2023), who instead find a weak positive correlation between the blackbody radius at peak optical light and the SMBH mass.
Alternately, the stellar debris may circularize rapidly without immediately accreting. Instead, the circularized debris may initially form a hot quasi-spherical pressure-supported envelope (Steinberg & Stone, 2024; Price et al., 2024), delaying rapid accretion onto the SMBH until the envelope cools and contracts to form a disk (Loeb & Ulmer, 1997; Coughlin & Begelman, 2014; Metzger, 2022). If the pressure-supported envelope model is correct, the timing of the super-Eddington accretion phase may be delayed by several hundred days relative to the MOSFiT models presented in Section 3.2, suggesting that it may still be possible for super-Eddington accretion to power the delayed radio emission observed in some TDEs even if the radio onset time occurs after MOSFiT’s nominal $\dot{M}\sim 0.3\;\dot{M_{\rm Edd}}$ transition point rather than prior to it (Metzger, 2022; Sarin & Metzger, 2024). In this case, the peak of the X-ray light curve will also be delayed relative to peak optical light and may provide a better estimate of the peak accretion time (Metzger, 2022). This is a likely explanation for the delayed X-rays seen in ASASSN-15oi (Figure 2 and Hajela et al. 2024). In scenarios where the early TDE UV/optical emission instead originates from a quasi-spherical envelope (Metzger, 2022), the initial size of the envelope (and hence its photosphere radius) increases with the SMBH mass; however, the cooling timescale over which the envelope contracts is also longer for lower SMBH mass, such that in the simplest scenario of passive cooling, the photosphere radius at a fixed time after the disruption is roughly independent of the SMBH mass. Given the low significance of our possible correlation between SMBH mass and delayed radio emission, this scenario is also consistent with our results.
4.2.2 Relativistic jets
Super-Eddington accretion can also power radio-luminous relativistic jets (De Colle et al., 2012). The two TDEs that show the most dramatic late-time radio brightenings (AT2018hyz and ASASSN-15oi) are both inconsistent with the Fundamental Plane at late times. This is clearest for ASASSN-15oi (stars in Figure 5), which shows two distinct radio emission episodes (Figure 2). ASASSN-15oi’s emission is consistent with the Fundamental Plane prior to and during the first radio flare, but moves significantly further away from the plane during the second flare, as the X-rays fade at the same time that the radio dramatically brightens. This suggests two different physical origins for the two radio flares, as proposed by Hajela et al. (2024). Several other TDEs in our sample possibly show a similar evolution away from the Fundamental Plane at late times (AT2019ehz, PS16dtm, and ASASSN-19bt), which could be confirmed by continued late-time radio monitoring accompanied by deeper X-ray observations.
The rapid radio brightening seen in these events may be powered by a jet (or wind) launched during a period of high accretion onto their SMBHs, when they are not expected to be on the Fundamental Plane. This is broadly consistent with the MOSFiT accretion rate predictions at the outflow launch times for ASASSN-15oi, PS16dtm, and AT2019ehz shown in Figures 2 and 3, although the systematic uncertainties on the SMBH masses for all three TDEs are large. The launch time for AT2018hyz shown in Figure 2 is taken from the non-relativistic outflow modeling by Cendes et al. (2022b) and does not coincide with MOSFiT’s prediction for super-Eddington accretion. This may either suggest that MOSFiT places the time of peak accretion too early, or that AT2018hyz’s radio emission is powered by an off-axis jet launched much earlier than the Cendes et al. (2022b) prediction but not visible until late times because of relativistic beaming (Matsumoto & Piran, 2023) or delayed disk-black hole spin alignment (Teboul & Metzger, 2023; Lu et al., 2024). Further radio monitoring of AT2018hyz, including Very Long Baseline Interferometry observations, may distinguish among these possibilities.
We also note that for TDEs, which exhibit rapidly-evolving accretion rates rather than the steady-state accretion typically seen in AGN, simultaneous radio and X-ray observations may not always be probing the same accretion rate if the radio emission is significantly spatially separated from the X-rays. In AGN on the Fundamental Plane, the radio emission is typically dominated by flat-spectrum emission from the jet core (Falcke et al., 2004) and thus reflects the current (or recent) accretion rate, as does the X-ray emission (which is presumed to originate close to the SMBH in both AGN and TDEs). In contrast, if the radio emission in our TDEs is dominated instead by optically-thin synchrotron emission from the forward shock between the ISM and the leading edge of an expanding jet (as first proposed for Swift J1644+57; e.g., Zauderer et al. 2011), then it may instead reflect the accretion rate at the time the jet was launched, which could be significantly different from the current accretion rate. This could explain why ASASSN-15oi and AT2018hyz deviate from the Fundamental Plane even at very late times. A similar argument has previously been suggested to explain why AGN don’t always follow the Fundamental Plane at low SMBH masses similar to those of TDE hosts (Gültekin et al., 2022).
4.2.3 Highly sub-Eddington accretion
Some delayed radio emission in TDEs may also be powered by highly sub-Eddington accretion. To assess this possibility, we must accurately determine the timing of the $\dot{M}\;\sim 0.03\dot{M}_{\rm Edd}$ state transition, when the disk is expected to transition from geometrically thin back to geometrically thick. In our Case 1 prompt accretion scenario, we can straightforwardly compute this timescale using our $\dot{M}$ curves from MOSFiT (Section 3.3). In Case 2 delayed accretion scenarios, the situation is less clear cut. If the cooling envelope model developed by Metzger (2022) is correct, then assuming that the envelope contracts to a disk faster than the mass fall-back reaches $\dot{M}\;\sim 0.03\dot{M}_{\rm Edd}$, and neglecting any spreading disk phase, the timescale estimates for this state transition that we obtain with MOSFiT while neglecting the envelope phase may still be accurate. [^16] We note that even if the MOSFiT prediction for the timing of the $\dot{M}\;\sim 0.03\dot{M}_{Edd}$ transition is underestimated, parameters such as $R_{\rm ph,0}$ that are primarily determined by the photometry around peak optical light are still reliable. We find that some TDEs in our sample (e.g. AT2018hco) do indeed exhibit delayed radio emission consistent with an outflow launched at approximately the same time that MOSFiT predicts this state transition will occur (Figure 2).
However, the existence of UV plateaus in many TDEs that remain well-fit by a thin disk model for years post-disruption (e.g., van Velzen et al. 2019a; Mummery & Balbus 2020; Mummery et al. 2024a, b) may suggest that in reality the $\dot{M}\;\sim 0.03\dot{M}_{Edd}$ transition is also delayed relative to the MOSFiT prediction, at least for some TDEs. [^17] This is not always the case; Mummery et al. (2024a) apply the thin disk model FitTeD to AT2019dsg and find that the onset of its second radio flare approximately coincides with the calculated $\dot{M}\;\sim 0.03\dot{M}_{Edd}$ state transition, just as MOSFiT does (Figure 3). Furthermore, in contrast to the hard X-ray spectra typical of AGN accreting at highly sub-Eddington rates, the X-ray emission from TDEs on timescales of years is often soft, consistent with thermal thin-disk emission extending to late times (e.g., Auchettl et al. 2017; Gezari et al. 2017; Wen et al. 2020; Jonker et al. 2020; Guolo et al. 2023 – but see also Komossa et al. 2004; Saxton et al. 2021 for discussion of X-ray selected TDEs that do show spectral hardening on timescales of years). We therefore must consider the full multiwavelength properties of each TDE in our sample to assess whether MOSFiT can provide an accurate estimate of its state transition timing, or whether other accretion probes (such as the X-ray emission) should be preferred.
The strongest evidence that some of our delayed radio TDEs may have entered the highly sub-Eddington phase comes from comparing their radio and X-ray emission. Many of our delayed radio TDEs (e.g., ASASSN-14ae, AT2018fyk, and AT2019teq) align with the Fundamental Plane at late times (Figure 5). This may indicate that these TDEs have indeed transitioned to an analog of the low-hard state seen in XRBs at late times and, like the AGN and XRBs used to calibrate the original Fundamental Plane relationship, are accreting at a small fraction of the Eddington rate at the time of observation. Indeed, AT2018fyk’s and AT2019teq’s X-ray spectra show significant hardening at late times, consistent with this picture (Wevers et al., 2019b, 2021; Yao & Guolo, 2022; Guolo et al., 2023; Pasham et al., 2024). This is also consistent with the radio non-detections of AT2018fyk at very early times when the X-rays were significantly brighter and softer, as outflows are not expected prior to this state transition when the accretion rate is higher but likely not super-Eddington given the high SMBH mass (Wevers et al., 2019b). ASASSN-14ae is the only TDE in our sample to show a correlated increase in both radio and X-ray emission at late times (Figure 2), but its late-time X-ray emission remains too faint for a detailed spectral analysis (Jonker et al., 2020). ASASSN-14ae is also the oldest TDE in our sample and has the longest delay between peak optical light and first radio detection. As illustrated by Figure 2, in order to explain ASASSN-14ae’s radio emission as an outflow launched by a state transition in the accretion flow at $\dot{M}\sim 0.03\dot{M}_{\rm Edd}$, accretion onto the SMBH must be significantly delayed relative to the $\dot{M}$ curve computed by MOSFiT (by $\sim 2000$ days). Thus, further monitoring of ASASSN-14ae’s radio and X-ray emission is necessary to ascertain the nature of the late radio rise in this TDE.
One caveat to this picture is that while some models predict a correlation between the between the SMBH mass and the timescale required to reach the low-hard state transition (typically hundreds of days; e.g., De Colle et al. 2012; Metzger 2022), we observe no clear correlation in our data between the SMBH mass and the onset or peak timescales of the radio emission, or between the SMBH mass and the peak radio luminosity. However, as noted in Section 3.2 and C24, the exact onset time of radio emission is poorly constrained for some events. In particular, seven of our TDEs have no radio observations at $t<180$ d and others are only sparsely observed on these timescales, meaning that their prompt radio emission could have been missed entirely. If these TDEs were misclassified in our analysis, then the simple statistical tests we conduct here may not accurately reflect the underlying population. We therefore do not rule out these models immediately based on the lack of an obvious trend between SMBH mass and radio onset time, but instead choose to defer a more in-depth analysis to when a larger sample of TDEs with well-sampled radio light curves at both early and late times is available.
4.3 Origin of the prompt radio emission
Finally, we briefly consider the origin of prompt radio emission in our TDEs. While various processes can delay accretion onto the SMBH relative to the $\dot{M}$ curves predicted by MOSFiT (e.g., Metzger 2022), it is more difficult to devise a mechanism by which the SMBH accretion could peak earlier than the MOSFiT prediction. This suggests that some other mechanism occurring earlier in the disruption process may be required to explain radio emission in some TDEs at early times – perhaps collision-induced outflows launched during the initial circularization of the stellar debris stream (Dai et al., 2013; Guillochon & Ramirez-Ruiz, 2015; Jiang et al., 2016; Sądowski et al., 2016; Lu & Bonnerot, 2020), or the interaction of the unbound portion of the stellar debris stream with the circumnuclear medium (Krolik et al., 2016; Yalinewich et al., 2019). However, in the latter model we would not expect any correlation with the Fundamental Plane, in contrast to our observations (Figure 5).
The observed adherence of our prompt radio TDEs to the Fundamental Plane may instead indicate that a small amount of stellar material does manage to accrete onto the SMBH even at early times. However, the X-ray emission of most Prompt radio TDEs, including the canonical event ASASSN-14li (squares in Figure 5), is much softer than that of the AGN that define the Fundamental Plane, arguing against a standard XRB-like low-hard state for this early phase (Pasham & van Velzen, 2018; Guolo et al., 2023). Our other highlighted Prompt radio TDE, AT2019azh, has one of the most extensive radio and X-ray datasets of any TDE; the data included in Figure 5 span $\sim 40-800$ days post-discovery (blue-edged diamonds). Interestingly, it remains roughly consistent with the Fundamental Plane for this entire time, despite late-time changes in its X-ray hardness that suggest possible XRB-like state changes in its accretion flow (Sfaradi et al., 2022). The proximity of the Prompt radio TDEs to the Fundamental Plane could conceivably be a coincidence given the small sample size, as noted by Pasham & van Velzen (2018). Indeed, the X-ray non-detections of AT2020vwl, AT2019qiz (at early times), and ASASSN-19bt (at late times) are potentially consistent with deviations from the Fundamental Plane, but deeper observations would be required to confirm this. As the number of TDEs in the published literature with sufficiently sensitive high-cadence radio and X-ray detections at early times remains extremely small, further exploration of these possibilities will require a larger sample of promptly-observed radio TDEs and is beyond the scope of this work.
5 Conclusions
We present an in-depth multi-wavelength analysis of a sample of 31 optical TDEs exhibiting radio emission on a variety of timescales, with the goal of elucidating the process(es) responsible for delayed radio emission in some events. We find that the multi-wavelength properties of our sample are as diverse as their radio properties, but nevertheless accretion-powered outflows remain an attractive option to explain delayed low-luminosity radio emission in many TDEs. Our primary conclusions are the following:
  • We use the TDE modeling code MOSFiT to predict the accretion rate onto the SMBH as a function of time for each TDE and estimate the timing of likely state transitions in the accretion flow. For a few TDEs (e.g., AT2018hco, the second radio flare in AT2019dsg), the transition to a highly sub-Eddington state $\dot{M}<0.03\dot{M}_{Edd}$ is roughly coincident with the launch time inferred for the delayed radio outflow, possibly suggesting that their radio emission is related to this state transition.
  • A larger number of TDEs may have radio outflows launched coincident with the $\dot{M}\sim 0.03\dot{M}_{Edd}$ state transition if this transition is delayed by several hundred days relative to the MOSFiT prediction. This is also supported by the X-ray behavior of these TDEs, which are consistent with the Fundamental Plane of black hole activity at late times. For ASASSN-14ae, this transition may happen as late as $\sim 2000$ days post-discovery.
  • Other TDEs (e.g., ASASSN-15oi, AT2018hyz) are much more radio-bright/X-ray dim than is typical for black holes accreting at a few percent of the Eddington rate, as seen by their increasing distance off the BH fundamental plane with time (Figure 5). We suggest that this sub-group of Delayed radio TDEs is powered instead by jets or outflows launched during a period of super-Eddington accretion. For these TDEs, the lack of bright radio emission at early times then implies either that the radio emission is powered by a promptly-launched off-axis jet initially suppressed by relativistic beaming, or that super-Eddington accretion is delayed by several hundred days relative to peak optical light.
  • TDEs with delayed radio emission may be identifiable at early times by their larger optical photospheric radii and lack of He emission at peak. If this indicates slower debris circularization and delayed-onset accretion as previously suggested by (Nicholl et al., 2022), the super-Eddington accretion phase would also be later than MOSFiT’s prediction and some delayed radio TDEs could be powered by non-relativistic outflows launched during this phase. If this correlation between TDE optical and radio properties is confirmed in a larger sample of TDEs, it may provide an important tool to predict which TDEs are likely to brighten in the radio at late times, enabling more efficient use of limited follow up resources.
  • Taken together, our analyses suggest that in TDEs the accretion rate commonly peaks several hundred days after peak optical light and then declines, driving radio-bright outflows in both the super-Eddington and highly sub-Eddington regimes. This may support delayed circularization models for the optical emission or the recent cooling envelope model proposed by Metzger (2022), in which the debris circularizes promptly but does not actually accrete until it has sufficiently cooled.
  • Finally, while a definitive assessment is precluded by the scarcity of available data, the onset of prompt radio emission in some TDEs may be too early to be plausibly explained by accretion onto the SMBH, suggesting that other phases of the disruption process may also be capable of powering fast outflows.
  • Our results strongly motivate systematic, consistent radio, optical, and X-ray monitoring campaigns of TDEs, beginning as soon as possible after discovery and extending to years post-disruption. Rapid public classification of new TDEs is also essential to ensure adequate multi-wavelength coverage at both early and late times, which is the only way to properly characterize the rates and origins of prompt and delayed radio emission in TDEs. Such campaigns are underway and will be reported in future work.
As larger samples of TDEs are discovered and the time cost of intensive targeted follow up becomes prohibitive, all-sky radio surveys will play an increasingly important role in systematically monitoring the full TDE population. Our independent VLASS discovery of the dramatic rebrightening of ASASSN-15oi (Hajela et al., 2024) demonstrates the value of unbiased radio surveys for uncovering unexpected behavior – it is unlikely that late-rising radio emission in TDEs would have been explored at all, without the initial VLASS detection of this TDE. Our few VLASS detections (Appendix A) confirm that $2-6$ years post-disruption may be a particularly fruitful time to search for late-rising low-frequency radio emission in optically selected TDEs.
All-sky radio surveys are now also uncovering new nuclear transients that may be TDEs missed at other wavelengths (e.g., Anderson et al. 2020; Ravi et al. 2022; Somalwar et al. 2023b, a; Dykaar et al. 2024). Our results suggest that deep late-time X-ray observations of these ambiguous radio transients may be one way to gain insight even when optical and X-ray observations extending to several years prior to the radio discovery are not available. While VLASS is the deepest radio survey of the northern sky conducted to date, many of the TDE outflows observed in the targeted radio observations of our sample are too faint to be detected in VLASS for all but the nearest TDEs, suggesting that VLASS’s limited sensitivity, frequency coverage, and time sampling may still be inadequate to fully characterize this previously unappreciated late phase of TDEs’ radio evolution. The planned ngVLA combined with wide-field radio surveys with the SKA will greatly increase our sensitivity to both low-luminosity radio outflows from TDEs and to distant jetted TDEs (Donnarumma et al., 2015; van Velzen et al., 2018), providing critical support for the large numbers of higher-redshift TDEs expected from Rubin Observatory and other upcoming optical and multi-wavelength surveys.
We thank Sjoert van Velzen, Adelle Goodwin, James Miller-Jones, and Nayana AJ for useful conversations. The National Radio Astronomy Observatory (NRAO) is a facility of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the CIRADA cutout service at URL cutouts.cirada.ca, operated by the Canadian Initiative for Radio Astronomy Data Analysis (CIRADA). CIRADA is funded by a grant from the Canada Foundation for Innovation 2017 Innovation Fund (Project 35999), as well as by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec, in collaboration with the National Research Council of Canada, the US National Radio Astronomy Observatory and Australia’s Commonwealth Scientific and Industrial Research Organisation. In addition to Swift observations proposed by us, we have also used the Swift archive at https://swift.gsfc.nasa.gov/archive/. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. KDA and CTC acknowledge support provided by the NSF through award SOSPA9-007 from the NRAO and award AST-2307668. KDA gratefully acknowledges support from the Alfred P. Sloan Foundation. NF acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2137419. BDM was supported in part by the National Science Foundation (grant No. AST-2009255). The Flatiron Institute is supported by the Simons Foundation. MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. ER-R was supported in part by the Heising-Simons Foundation, NSF (AST-2150255 and AST-2307710), Swift (80NSSC21K1409,80NSSC19K1391) and Chandra (22-0142). FDC acknowledges support from the UNAM-PAPIIT grant IN113424. This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). RC and RM acknowledge the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, for its hospitality while this work was completed. , CASA (McMullin et al., 2007), pwkit (Williams et al., 2017)
Appendix A VLASS search for late-rising radio emission from known TDEs
The advent of the VLASS, the most sensitive radio survey of the full northern sky to date, provides an exciting new opportunity to search for late-rising radio emission from known TDE candidates. VLASS began on 2017 September 7 and is using the VLA to map 80% of the radio sky in three epochs spanning seven years (Lacy et al., 2020). All VLASS observations are taken at $2-4$ GHz (S-band) and the typical rms sensitivity of the single-epoch images is $\sim 100$ $\mu$Jy.
Given the newly-recognized prevalence of long-lasting radio emission in TDEs, all-sky radio surveys also open up a complementary method of discovering new TDEs in the nearby Universe. Blind radio searches with VLASS and radio surveys with the SKA precursors are now finding dozens of nuclear transients that may be TDEs, but interpretation is difficult without multi-frequency radio follow up (e.g., Anderson et al. 2020; Ravi et al. 2022; Somalwar et al. 2023b, a; Dykaar et al. 2024). The rates and higher average radio luminosities of transients discovered in these blind searches suggest that not all of them are drawn from the same underlying population as the TDEs studied in this work, but it is not clear if this reflects a true dichotomy or a bias in host environment – for example, TDEs in dense gas and dust-rich environments may produce stronger radio emission (Generozov et al., 2017), but may have weaker observed optical and X-ray emission due to extinction. VLASS allows us to easily compare the radio properties of optical TDEs to those of other nuclear transients.
Here, we use VLASS to investigate if the late-time radio properties of the optical TDEs studied in this work are representative of the broader, heterogeneously-selected full known TDE population, which includes TDEs discovered via a wide range of methods (e.g., as flares in all-sky X-ray or infrared surveys). We compiled a broad list of TDE candidates from the literature, consisting of objects from the Open TDE Catalog [^18] The Open TDE Catalog, https://github.com/astrocatalogs/tidaldisruptions, provides a largely-complete list of TDEs reported in the literature through mid-2018. It includes some objects with ambiguous classification as well as transients widely accepted by the community as TDEs. (Guillochon et al., 2017) and the TDE sample from French et al. (2020) that are within the VLASS footprint. Our target list is optimized for completeness rather than purity; we do not make any cuts based on TDE redshift, age, or discovery method. Our final sample contains 91 objects discovered between 1990 and 2019, including 40 optical TDE candidates, 39 X-ray TDE candidates, and 18 TDE candidates identified via other methods, including nuclear infrared flares and bright transient gamma-ray emission (Figure 6). The redshifts of our TDE candidates range from 0.000271 to 1.1853 (not including one X-ray transient in the Chandra Deep Field South with a photometric redshift $z\sim 2$; Bauer et al. 2017); the median redshift of the sample is $z=0.088$. Approximately two thirds of the TDE candidates in this list had not received any published targeted radio follow up at the time of our search.
We cross-matched the locations of these 91 TDE candidates with our Source Extractor catalog of possible VLASS source detections, publicly available here: 10.5281/zenodo.4895112. The creation of this catalog is described in Stroh et al. (2021). For this search, we utilized the Epoch 1 quick-look images provided by NRAO. [^19] For additional information about the VLASS quick-look images and their limitations, see https://science.nrao.edu/vlass/data-access/vlass-epoch-1-quick-look-users-guide. The first half of the data (VLASS1.1) was collected between September 2017 and February 2018 and suffers from additional uncertainties in the absolute flux calibration scale; the peak flux densities of sources fainter than 1 Jy are systematically low by 15%, with an additional systematic scatter of ±$\sim\pm 8$%. The second half of the data (VLASS1.2) was collected between February and July 2019; the total flux densities of faint sources in these data are low by $\sim 8$% with a scatter of $\sim 5\%$. As our primary goal is simply to determine the possible presence of transient radio emission, these calibration uncertainties are acceptable for our purposes.
We identified 5 TDEs located within 5 arcsec of a VLASS source detected with a signal-to-noise ratio $>5$ in our catalog, which are presented in Table 3. We use the imfit command within CASA to fit an elliptical Gaussian to each TDE and take the peak brightness to be equal to the flux density, as expected for an unresolved point source. Interestingly, each of these TDEs was discovered via a different method: NGC 5905 was one of the first TDE candidates ever reported and was identified as an X-ray flare from the center of a non-AGN galaxy (Bade et al., 1996); Swift J1644+57 was discovered as a long-lasting, luminous gamma-ray and X-ray transient (Bloom et al., 2011; Burrows et al., 2011; Levan et al., 2011); ASASSN-15oi was discovered in an optical survey and classified as a TDE based on its nuclear location, photometric evolution, and spectroscopic properties (Holoien et al., 2016a); J100933 was discovered as a mid-infrared flare (Wang et al., 2018); and F01004-2237 was identified based on changes to its optical spectrum observed serendipitously during a survey of nearby ultra-luminous infrared galaxies (Tadhunter et al., 2017). Two of these five objects, Swift J1644+57 and ASASSN-15oi, have well-studied radio counterparts previously reported in the literature (Zauderer et al., 2011; Berger et al., 2012; Wiersema et al., 2012; Zauderer et al., 2013; Yang et al., 2016; Eftekhari et al., 2018; Cendes et al., 2021b; Horesh et al., 2021a; Hajela et al., 2024). Swift J1644+57’s radio emission, first detected within days of its discovery, is well-explained as synchrotron emission from an initially relativistic jet viewed on-axis (or slightly off-axis; see Beniamini et al. 2023). Its detection in VLASS epochs 1 and 2 was previously reported by Cendes et al. (2021b). The epoch 1 VLASS detection of ASASSN-15oi, also reported by Horesh et al. (2021a) and Hajela et al. (2024), was the first evidence for a dramatic second brightening of its radio emission, providing the initial impetus for the deeper late-time observations of other TDEs presented in C24 and this work. NGC 5905 was the first TDE with a follow-up radio observation to search for jet emission. A VLA observation $\sim$ 6 years after the X-ray peak provided an upper limit of $<0.15$ mJy at 8.46 GHz for a central point source (Komossa & Dahlem, 2001). It was also observed as part of the NRAO VLA Sky Survey (NVSS). The NVSS detection of extended radio emission of NGC 5905 (21 mJy at 1.4 GHz) was found to be consistent with the starburst properties of the host galaxy and consistent with pre-outburst extended radio emission (Komossa, 2002; Raichur et al., 2015). The VLASS is less sensitive to extended emission due to the survey’s higher resolution, but the VLASS flux densities of NGC 5905 are approximately constant in time and are also likely consistent with unrelated emission from the host galaxy. We include NGC 5905’s VLASS flux density in Table 3 for completeness but do not investigate this source further. We also choose not to further observe F01004-2237, as the nature of this source is disputed. It was initially proposed to be a TDE (Tadhunter et al., 2017), later re-classified as an unusual AGN Bowen fluorescence flare (Trakhtenbrot et al., 2019), and finally recently suggested to be a repeating partial TDE (or two independent TDEs) after a second optical flare was observed in September 2021 (Sun et al., 2024). The continued activity of this source motivates future multi-wavelength monitoring to distinguish between the AGN and repeated or partial TDE scenarios.
We triggered additional multi-frequency follow up observations of the two most promising radio TDE candidates in our sample, ASASSN-15oi and J100933. We observed these targets with the VLA on 2020 May 20 (program 20A-492, PI: Alexander) and 2021 June 10 (program 21A-303, PI: Hajela). Observations were conducted at $1-12$ GHz (L, S, C, and X bands) when the VLA was in its C configuration. The data were reduced in CASA using standard procedures and flux densities were measured using the imtool fitsrc command within the pwkit package (Williams et al., 2017). We have verified that this gives similar results to measuring the flux densities with imfit. The follow up observations of ASASSN-15oi are presented in Hajela et al. (2024). We include the measured flux densities of J100933 in Table 3 together with their statistical errors. The systematic uncertainty on the flux calibration for these observations is 5% at all frequencies. In our 2021 observations, J100933 exhibits a rising spectrum with a peak around 9 GHz. We find that the flux density of J100933 at 3 GHz increased slightly between the time of the VLASS epoch 1 detection and our follow up observations two years later and then remained approximately constant for the next three years. Given the additional flux calibration uncertainties in the quick look images for epoch 1 of VLASS, this is roughly consistent with no significant temporal evolution over the full 5-year period. We thus conclude that the radio emission in this target is likely unrelated to the transient, although longer-term multi-frequency monitoring may be needed to confirm this due to the age of the TDE ($>11$ yr).
We also visually inspected the coordinates of each of the 91 TDEs in our sample in VLASS epochs 2.1, 2.2, and 3.1 quicklook data using the CIRADA cutout server to see if any of them appeared in the radio after VLASS epoch 1. We find only one new TDE undetected in VLASS epoch 1 that “turns on” in subsequent data: AT2018hyz, which is not detected in epochs 1 or 2 but appears as a bright point source in VLASS epoch 3. The VLASS data for AT2018hyz have already been presented in Cendes et al. (2022b) and C24; we therefore do not repeat them here. We also continue to detect the five sources visible in our Epoch 1 search. NGC 5905, F01004-2237, and J100933 exhibit roughly constant radio brightness across all VLASS epochs, while Sw J1644+57 and ASASSN-15oi slowly fade. We include the flux densities of these five TDEs in all available VLASS epochs in Table 3.
In summary, our VLASS search of 91 TDE candidates resulted in three detections of transient radio emission associated with genuine TDEs (two visible in all three VLASS epochs, and one that is only detected in VLASS epoch 3), two radio detections of TDE host galaxies, and one radio detection of a likely non-TDE transient (F01004-2237) or its host galaxy. Although our statistical constraining power is limited due to the small number of detections, the diverse multi-wavelength properties of these objects suggest no clear preference for optically selected TDEs among the TDE sub-population that is radio-bright at late times, at least at the bright end of the radio luminosity function. The low detection fraction seen in our search is consistent with early searches for late-time radio emission in TDEs, which concluded that powerful relativistic jets like that seen in the TDE Swift J1644+57 are rare (Komossa, 2002; van Velzen et al., 2013). Due to the lower sensitivity of available radio telescopes at the time and the relatively large distances of many early TDE candidates, these early searches were not sensitive to lower-luminosity radio emission similar to that seen in more recently-discovered nearby TDEs, many of which have peak radio luminosities $\nu L_{\nu}\lesssim 10^{38}$ erg s-1 (Alexander et al. 2020; C24). Even our VLASS search can only detect $\nu L_{\nu}=10^{38}$ erg s-1 outflows out to $z\sim 0.065$, meaning that such emission would be missed in $\gtrsim 60$% of our sample even if all TDEs were captured at or near their peak radio brightness (which is unlikely due to the broad distribution of TDE ages in our sample). Our results suggest that $\sim 2-6$ years post-discovery may be a particularly fruitful timescale to search for late-rising radio emission in TDEs at GHz frequencies and demonstrates the value of long-duration monitoring, which will be enabled by future generations of more sensitive radio surveys.
Figure 6: The ages of the TDEs in our VLASS sample at the time each was observed in VLASS epoch 1. We detect radio emission from 5 TDEs out of a total sample of 91. Three detections are likely host galaxy emission (NGC 5905, F01004-2237, and J100933), while two are confirmed to be transient radio emission associated with the TDE (Sw J1644+57 and ASASSN-15oi). Both confirmed TDE radio detections occur on timescales of $2-6$ years post-discovery (rest frame), suggesting that this may be an interesting timescale for deeper follow up. (A sixth TDE, AT2018hyz, is detected in VLASS epoch 3 at a similar timescale post-disruption.)
Table 3: Radio observations of TDE candidates first detected in Epoch 1 of VLASS. We show both statistical errors from our source fitting procedure and systematic errors from the absolute flux density calibration scale of the data. Note that the VLASS errors are asymmetric to reflect that the peak flux densities of the quick look images are on average biased low relative to pointed observations (see text for details).
TDE Name Redshift Earliest Detection Observation Program Frequency Flux density Notes
(any wavelength) Date (GHz) (mJy $\pm$ stat $\pm$ sys)
NGC 5905* 0.01124 1990 Jul 11 2017 Nov 11 VLASS 1.1 3.0 $1.59\pm 0.12^{+0.08}_{-0.37}$ X-ray TDE
2020 Aug 3 VLASS 2.1 3.0 $1.96\pm 0.16^{+0.10}_{-0.45}$
2023 Feb 9 VLASS 3.1 3.0 $1.65\pm 0.12^{+0.08}_{-0.38}$
F01004-2237* 0.117835 2010 Jun 25 2018 Feb 13 VLASS 1.1 3.0 $3.47\pm 0.13^{+0.17}_{-0.80}$ AGN flare
2020 Nov 1 VLASS 2.1 3.0 $3.39\pm 0.15^{+0.17}_{-0.78}$
2023 Jun 12 VLASS 3.1 3.0 $4.1\pm 0.2^{+0.2}_{-0.9}$
Swift J1644+57 0.3543 2011 Mar 25 2017 Oct 7 VLASS 1.1 3.0 $1.32\pm 0.11^{+0.07\dagger}_{-0.30}$ gamma-ray TDE
2020 Aug 4 VLASS 2.1 3.0 $0.64\pm 0.16_{-0.15}^{+0.03\dagger}$
2023 Feb 11 VLASS 3.1 3.0 $0.346\pm 0.08^{+0.02}_{-0.08}$
ASASSN-15oi 0.02 2015 Aug 14 2019 Jul 1 VLASS 1.2 3.0 $9.0\pm 0.2_{-1.0}^{+0.6\ddagger}$ optical TDE
2022 Feb 15 VLASS 2.2 3.0 $4.95\pm 0.14^{+0.35}_{-0.54}$
J100933* 0.0719 2009 Feb 20 2019 Apr 18 VLASS 1.2 3.0 $1.88\pm 0.15^{+0.13}_{-0.21}$ mid-IR flare
2021 Jun 10 21A-303 1.263 $1.59\pm 0.16\pm 0.08$
2021 Jun 10 21A-303 1.775 $2.15\pm 0.10\pm 0.11$
2021 Jun 10 21A-303 2.5 $2.49\pm 0.04\pm 0.12$
2021 Jun 10 21A-303 3.5 $3.04\pm 0.03\pm 0.15$
2021 Jun 10 21A-303 3.0 $2.85\pm 0.04\pm 0.14$ full S band image to match
VLASS (combines above 2 rows)
2021 Jun 10 21A-303 5.0 $3.65\pm 0.03\pm 0.18$
2021 Jun 10 21A-303 7.0 $4.12\pm 0.04\pm 0.21$
2021 Jun 10 21A-303 9.0 $4.20\pm 0.03\pm 0.21$
2021 Jun 10 21A-303 11.0 $4.06\pm 0.04\pm 0.20$
2021 Dec 1 VLASS 2.2 3.0 $2.47\pm 0.11^{+0.17}_{-0.27}$
2024 Apr 30 VLASS 3.2 3.0 $2.46\pm 0.13^{+0.17}_{-0.27}$
*indicates origin of radio emission is either ambiguous or unrelated to the TDE (see text).
Consistent with values previously reported in Cendes et al. (2021b).
Consistent with value previously reported in Hajela et al. (2024).
Appendix B Data Tables
In this section, we present a table listing the original references for all published optical/UV, X-ray, and radio data used in our analysis (Table 4) and complete data tables for all new observations. Our new radio observations of the TDE AT2019qiz (Section 2.1) are listed in Table 5 and new Swift-XRT observations of our sample (Section 2.2) are listed in Table 6. Table 6 also includes public archival Swift observations of our TDEs that have not previously been published, which were reduced using the same method as our new observations (Section 2.2). We will also make the full multi-wavelength dataset for each TDE in our sample publicly available online through OTTER (Franz et al in prep).
Our radio data first published in C24 were taken with the VLA (programs 20B-492, PI: Alexander; 21-303, PI: Hajela; 22B-205, PI: Cendes), MeerKAT (programs SCI-20210212-YC-01, DDT-20220414-YC-01, and SCI-20220822-YC-01; PI: Cendes), and the Australia Telescope Compact Array (ATCA; programs C3472, PI: Cendes and C3325, PI: Alexander). The full list of observations, measured radio flux densities and upper limits, and details of the data reduction for these 24 objects are given in Cendes et al. (2022b), C24, and references therein. Our new radio observations of AT2019qiz are described in Section 2.1; our measured flux densities and the observing program used to collect each data point are listed in Table 5.
Sixteen of our TDEs were included in the uniform X-ray analysis published by Guolo et al. (2023), which includes observations taken by XMM-Newton and by the X-ray Telescope (XRT, Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004). We use the data from Guolo et al. (2023) for 14 of these objects; for ASASSN-15oi, we instead use the more comprehensive dataset presented by Hajela et al. (2024), and for AT2019qiz, we use the X-ray dataset from Nicholl et al. (2024), which corrects the XRT photometry for the presence of a nearby background source unrelated to the TDE. We also include published observations from the Chandra X-ray Observatory for a few TDEs: ASASSN-14ae (Jonker et al., 2020), PS16dtm (Petrushevska et al., 2023), AT2017eqx (Nicholl et al., 2019), and AT2018hyz (Cendes et al., 2022b). For AT2018fyk, we also include the late-time X-ray dataset from Pasham et al. (2024) consisting of XRT, XMM, and Chandra observations at $t>1000$ d post-discovery. For ASASSN-19bt, we report XMM-Newton photometry from Holoien et al. (2019a) and Swift-XRT and Chandra photometry from Christy et al. (2024).
We include all of the optical/UV photometry from Nicholl et al. (2022) for the TDEs presented in that work (Table 4 also lists the original references for these data). For AT2018lna, AT2019teq, AT2020mot, AT2020opy, AT2020pj, and AT2020wey we include $g$ and $r$-band photometry from the Zwicky Transient Facility (ZTF; Bellm et al. 2019b) obtained from the Automatic Learning for the Rapid Classification of Events (ALeRCE) broker (Förster et al., 2021) and UVOT photometry from Hammerstein et al. (2023a). For AT2020neh, AT2020nov, and AT2020vwl, we include only $g$ and $r$-band photometry from ZTF from ALeRCE. For OGLE17aaj we include I-band photometry from Optical Gravitational Lensing Experiment (OGLE; Wyrzykowski et al. 2014) up until MJD = 58200. We include the UVOT photometry from Wevers et al. (2019b) for AT2018fyk and from van Velzen et al. (2021) for AT2018bsi. For ASASSN-19bt, we include both published photometry and the public light curve from the ASAS-SN Sky Patrol (Hart et al., 2023). Six TDEs in our sample (ASASSN-14li, ASASSN-15oi, AT2018hyz, ASASSN-19bt, AT2019teq, and AT2019ehz) show late-time plateaus in their optical and UV light curves that are not well-captured by MOSFiT, as discussed in Section 3.2. We therefore only use the early photometry from these events for our modeling. The cutoff date for each of these TDEs is listed in Table 4. Similarly, for PS16dtm we include only the photometry from Blanchard et al. (2017) as the additional photometry published by Petrushevska et al. (2023) is not well-captured by MOSFiT.
[…]
Table 4: References for all data used in our analysis. All data for which the reference is “This work” are given in Tables 5 and 6.
TDE Name References for Optical/ References for References for
UV photometry; time range X-ray data radio data
ASASSN-14ae Holoien et al. (2014); Jonker et al. (2020); C24
Hinkle et al. (2021); This work
Nicholl et al. (2022)
DES14C1kia None This work C24
ASASSN-14li Jose et al. (2014); Guolo et al. (2023); Alexander et al. (2016)
Holoien et al. (2016b); This work
Hinkle et al. (2021);
Nicholl et al. (2022);
$t\leq 57300$ MJD
iPTF15af Blagorodnova et al. (2019); This work Blagorodnova et al. (2019);
Nicholl et al. (2022) C24
ASASSN-15oi Holoien et al. (2016a); Hajela et al. (2024) Horesh et al. (2021a);
Hinkle et al. (2021); Hajela et al. (2024)
Nicholl et al. (2022);
$t\leq 57400$ MJD
iPTF16axa Hung et al. (2017); This work C24
Hinkle et al. (2021);
Nicholl et al. (2022)
PS16dtm Blanchard et al. (2017) Petrushevska et al. (2023); Blanchard et al. (2017);
This work C24
iPTF16fnl Blagorodnova et al. (2017); This work Blagorodnova et al. (2017);
Hinkle et al. (2021); Horesh et al. (2021b);
Nicholl et al. (2022) C24
OGLE17aaj Wyrzykowski et al. (2014); This work Stanway et al. (2017);
Gromadzki et al. (2019); C24
$t\leq 58200$ MJD
AT2017eqx Nicholl et al. (2019); Nicholl et al. (2019) (Chandra); Nicholl et al. (2019);
Hinkle et al. (2021); This work (XRT) C24
Nicholl et al. (2022)
AT2018zr van Velzen et al. (2019b); Guolo et al. (2023) van Velzen et al. (2019b);
Holoien et al. (2019b); C24
van Velzen et al. (2021);
Hinkle et al. (2021);
Nicholl et al. (2022)
AT2018bsi van Velzen et al. (2021) Guolo et al. (2023); C24
This work
AT2018dyb Leloudas et al. (2019); This work Holoien et al. (2020);
Holoien et al. (2020); C24
Hinkle et al. (2021);
Nicholl et al. (2022)
AT2018fyk Wevers et al. (2019b) Guolo et al. (2023); Wevers et al. (2019b);
Pasham et al. (2024) Wevers et al. (2021);
C24;
Cendes et al. (2024b)
AT2018hco van Velzen et al. (2021); Guolo et al. (2023); C24
Hinkle et al. (2021); This work
Nicholl et al. (2022)
AT2018hyz Gomez et al. (2020); Cendes et al. (2022b); Horesh et al. (2018);
van Velzen et al. (2021); Guolo et al. (2023) Gomez et al. (2020);
Hinkle et al. (2021); Cendes et al. (2022b);
Nicholl et al. (2022); C24
$t\leq 58800$ MJD
AT2018lna van Velzen et al. (2021); Guolo et al. (2023); C24
Hinkle et al. (2021); This work
Nicholl et al. (2022);
Hammerstein et al. (2023a)
ASASSN-19bt ASAS-SN Sky Patrol; Holoien et al. (2019a); Christy et al. (2024)
Holoien et al. (2019a); Christy et al. (2024)
Hinkle et al. (2021);
Nicholl et al. (2022);
$t\leq 58650$ MJD
AT2019azh van Velzen et al. (2021); Guolo et al. (2023); Goodwin et al. (2022);
Hinkle et al. (2021); This work Sfaradi et al. (2022)
Nicholl et al. (2022)
AT2019dsg van Velzen et al. (2021); Guolo et al. (2023); Cendes et al. (2021a);
Cannizzaro et al. (2021); This work Stein et al. (2021);
Hinkle et al. (2021); C24
Nicholl et al. (2022)
AT2019ehz van Velzen et al. (2021); Guolo et al. (2023) C24
Hinkle et al. (2021);
Nicholl et al. (2022);
$t\leq 58800$ MJD
AT2019eve van Velzen et al. (2021) This work C24
Hinkle et al. (2021)
Nicholl et al. (2022)
AT2019qiz van Velzen et al. (2021); Nicholl et al. (2024) This work;
Hinkle et al. (2021); Franz et al. (in prep)
Nicholl et al. (2020);
Nicholl et al. (2022)
AT2019teq ZTF $g,r$ (ALERCE); Guolo et al. (2023); Cendes et al. (2022a);
Hammerstein et al. (2023a); This work C24
$t\leq 58850$ MJD
AT2020pj ZTF $g,r$ (ALERCE); Guolo et al. (2023); C24
Hammerstein et al. (2023a) This work
AT2020mot ZTF $g,r$ (ALERCE); This work Liodakis et al. (2023);
Hammerstein et al. (2023a) C24
AT2020neh ZTF $g,r$ (ALERCE) This work Angus et al. (2022);
C24
AT2020nov ZTF $g,r$ (ALERCE) This work C24
AT2020opy ZTF $g,r$ (ALERCE); This work Goodwin et al. (2023b)
Hammerstein et al. (2023a)
AT2020wey ZTF $g,r$ (ALERCE); Guolo et al. (2023); C24
Hammerstein et al. (2023a) This work
AT2020vwl ZTF $g,r$ (ALERCE) Guolo et al. (2023); Goodwin et al. (2023a)
This work
Table 5: New radio observations of AT2019qiz included in this work. We show both statistical-only errors from our fitting procedure and additional systematic errors from the VLA’s absolute flux calibration.
MJD Time since Flux density Frequency Program
Discovery (d) (mJy $\pm$ stat $\pm$ sys) (GHz)
58752.41169 6.91 $0.033\pm 0.008\pm 0.002$ 15.0 VLA/19A-013
58773.36964 27.87 $0.093\pm 0.012\pm 0.005$ 9.0 VLA/19A-013
58799.36227 53.86 $0.15\pm 0.04\pm 0.007$ 5.0 VLA/19A-013
58799.36227 53.86 $0.65\pm 0.03\pm 0.03$ 9.0 VLA/19A-013
58823.29204 77.79 $0.40\pm 0.04\pm 0.02$ 5.0 VLA/19A-013
58823.29204 77.79 $1.33\pm 0.02\pm 0.07$ 9.0 VLA/19A-013
58871.11659 125.62 $1.41\pm 0.03\pm 0.07$ 5.0 VLA/19A-013
58871.11659 125.62 $2.40\pm 0.02\pm 0.12$ 9.0 VLA/19A-013
58981.73736 236.24 $1.64\pm 0.05\pm 0.08$ 5.0 VLA/20A-372
58981.73736 236.24 $0.81\pm 0.10\pm 0.04$ 9.0 VLA/20A-372
59213.07616 467.58 $1.366\pm 0.015\pm 0.07$ 5.0 VLA/20A-372
59213.07616 467.58 $0.695\pm 0.013\pm 0.03$ 9.0 VLA/20A-372
59378.77265 633.27 $0.48\pm 0.06\pm 0.02$ 5.0 VLA/21A-303
59378.77265 633.27 $0.260\pm 0.017\pm 0.013$ 9.0 VLA/21A-303
59916 1170.50 $0.618\pm 0.015\pm 0.03$ 1.285 MKT-22085
60028.00797 1282.51 $0.7\pm 0.2\pm 0.02$ 3.0 VLASS 3.1
[…]
Table 6: New Swift-XRT data used in this work. This table includes both our new late-time TOO observations and public archival observations re-reduced by us. The date column gives the mean observation time for each data point, while the time since discovery includes the observation length or the full time range for stacked observations. The luminosity is calculated assuming a $\Gamma=2$ power-law spectrum, as discussed in Section 2.2. All upper limits are $3\sigma$ (Gaussian-equivalent).
TDE Name Date Time since 0.3-10 keV Count Rate 0.3-10 keV Luminosity
(MJD) discovery (d) ($10^{-3}$ counts s-1) ($10^{41}$ erg s-1)
ASASSN14ae 59590.754 $2908.2^{+0.2}_{-8.1}$ $1.3^{+0.6}_{-0.4}$ $3.1^{+1.3}_{-1.0}$
ASASSN14ae 56779.066 $97^{+660}_{-94}$ $<0.401$ $<0.955$
ASASSN14ae 59891.566 $3209.06^{+0.17}_{-0.17}$ $2.1^{+1.4}_{-1.0}$ $5^{+3}_{-2}$
ASASSN14ae 59861.832 $3179.32^{+0.02}_{-0.59}$ $<5.95$ $<14.2$
ASASSN14ae 60419.68 $3737.168^{+0.003}_{-0.003}$ $<16.5$ $<39.3$
ASASSN14ae 60502.559 $3820.0^{+0.2}_{-0.6}$ $<10.3$ $<24.6$
ASASSN14ae 60591.594 $3909.1^{+0.4}_{-0.5}$ $<4.51$ $<10.7$
DES14C1kia 57028.68 $56.7^{+0.5}_{-0.5}$ $<2.52$ $<96.8$
DES14C1kia 57053.332 $81^{+46}_{-10}$ $<0.579$ $<22.3$
DES14C1kia 57119.262 $148^{+8}_{-3}$ $<1.61$ $<62.1$
DES14C1kia 57433.223 $461.2^{+8.0}_{-0.3}$ $<3.91$ $<150$
DES14C1kia 59095.594 $2123.6^{+1.5}_{-1.5}$ $<2.92$ $<112$
DES14C1kia 59585.453 $2613.453^{+0.011}_{-0.007}$ $<7.07$ $<272$
DES14C1kia 60410.473 $3438.5^{+0.5}_{-0.5}$ $<2.81$ $<108$
ASASSN14li 59550.832 $2567.2^{+0.2}_{-0.7}$ $<3.81$ $<1.96$
ASASSN14li 59896.770 $2913.170^{+0.010}_{-0.010}$ $<10.8$ $<5.55$
ASASSN14li 59926.395 $2942.794^{+0.005}_{-0.014}$ $<9.4$ $<4.82$
ASASSN14li 59957.629 $2974.0^{+1.6}_{-1.6}$ $<3.84$ $<1.97$
ASASSN14li 60117.195 $3133.60^{+0.13}_{-0.01}$ $<6.61$ $<3.39$
ASASSN14li 60535.023 $3551.423^{+0.006}_{-0.002}$ $<23.5$ $<12$
ASASSN14li 60712.695 $3729.095^{+0.003}_{-0.003}$ $<20.5$ $<10.5$
iPTF15af 57059.859 $23^{+30}_{-11}$ $<0.614$ $<5.04$
iPTF15af 57104.273 $67^{+5}_{-5}$ $<2.21$ $<18.2$
iPTF15af 57176.855 $140^{+9}_{-9}$ $<1.4$ $<11.5$
iPTF15af 59582.012 $2544.712^{+0.009}_{-0.009}$ $<5.54$ $<45.5$
iPTF16axa 57575.492 $38^{+45}_{-29}$ $<0.451$ $<7.2$
iPTF16axa 59582.238 $2044.841^{+0.010}_{-0.010}$ $<5.63$ $<89.9$
iPTF16axa 60607.914 $3070.517^{+0.002}_{-8.6}$ $<8.51$ $<136$
iPTF16axa 60624.219 $3086.822^{+0.003}_{-9.9}$ $<8.22$ $<131$
iPTF16axa 60624.938 $3087.541^{+0.002}_{-0.002}$ $<29.4$ $<470$
PS16dtm 57874.484 $262^{+900}_{-240}$ $<0.135$ $<1.15$
PS16dtm 59018.629 $1410^{+750}_{-210}$ $<0.556$ $<4.74$
iPTF16fnl 57681.734 $52^{+1000}_{-51}$ $0.27^{+0.11}_{-0.11}$ $0.09^{+0.03}_{-0.03}$
iPTF16fnl 59724.562 $2095.16^{+0.67}_{-1.40}$ $<1.94$ $<0.624$
iPTF16fnl 59779.367 $2149.97^{+0.63}_{-0.16}$ $<8.08$ $<2.59$
iPTF16fnl 59809.574 $2180.17^{+0.07}_{-0.01}$ $<6.21$ $<1.99$
iPTF16fnl 60453.547 $2824.147^{+5}_{-0.007}$ $<1.42$ $<0.456$
iPTF16fnl 60463.934 $2834.5^{+0.9}_{-0.2}$ $<3.09$ $<0.991$
iPTF16fnl 60629.207 $3000^{+5}_{-5}$ $<3.52$ $<1.13$
iPTF16fnl 60635.332 $3005.93^{+0.04}_{-0.04}$ $<4.64$ $<1.49$
iPTF16fnl 60652.441 $3023^{+4}_{-4}$ $<2.13$ $<0.683$
iPTF16fnl 60666.285 $3036.89^{+0.08}_{-3.90}$ $<3.72$ $<1.19$
iPTF16fnl 60677.164 $3047.76^{+7.20}_{-0.02}$ $<5.42$ $<1.74$
iPTF16fnl 60694.668 $3065^{+4}_{-4}$ $<9.53$ $<3.06$
iPTF16fnl 60703.297 $3073.9^{+0.1}_{-0.1}$ $<4.8$ $<1.54$
OGLE17aaj 57850.949 $96^{+150}_{-74}$ $1.5^{+0.4}_{-0.4}$ $28^{+7}_{-7}$
OGLE17aaj 58286.012 $531.0^{+1.7}_{-1.2}$ $<8.32$ $<155$
OGLE17aaj 59587.258 $1832.3^{+0.1}_{-0.1}$ $<5.09$ $<94.7$
OGLE17aaj 60431.996 $2677.0^{+0.2}_{-0.2}$ $<5.91$ $<110$
AT2017eqx 58368.141 $460^{+500}_{-420}$ $<0.869$ $<14.1$
AT2017eqx 59586.363 $1674.803^{+0.2}_{-0.2}$ $<4.63$ $<75.3$
AT2017eqx 60417.230 $2505.7^{+4.1}_{-0.7}$ $<1.24$ $<20.2$
AT2017eqx 60430.508 $2518.948^{+0.0019}_{-0.012}$ $<8.77$ $<143$
AT2017eqx 60471.750 $2560.190^{+0.003}_{-0.003}$ $<18.1$ $<295$
AT2018zr 60453.875 $2274.6^{+0.5}_{-0.6}$ $3.0^{+2.1}_{-1.5}$ $20^{+14}_{-10}$
AT2018zr 60557.391 $2378^{+1}_{-2}$ $2.8^{+1.2}_{-1.0}$ $19^{+8}_{-6}$
AT2018zr 60609.387 $2430.062^{+0.47}_{-0.003}$ $<5.19$ $<34.1$
AT2018bsi 58486.047 $269^{+57}_{-850}$ $<1.34$ $<4.42$
AT2018bsi 59586.457 $1369.245^{+0.010}_{-0.010}$ $<5.18$ $<17.1$
AT2018bsi 60457.453 $2240.2^{+2.3}_{-0.4}$ $<6.81$ $<22.4$
AT2018bsi 60553.332 $2336.12^{+1.4}_{-0.01}$ $<2.22$ $<7.33$
AT2018dyb 58459.211 $149^{+620}_{-140}$ $0.49^{+0.15}_{-0.15}$ $0.19^{+0.06}_{-0.06}$
AT2018dyb 59591.625 $1281^{+4}_{-6}$ $<1.78$ $<0.696$
AT2018dyb 59806.180 $1495.95^{+0.06}_{-0.03}$ $<10.6$ $<4.16$
AT2018dyb 59826.746 $1516.52^{+0.12}_{-0.03}$ $<7.66$ $<3$
AT2018dyb 60481.043 $2171^{+5}_{-5}$ $<5.34$ $<2.09$
AT2018dyb 60493.578 $2183.348^{+0.001}_{-0.008}$ $<16.8$ $<6.56$
AT2018dyb 60618.914 $2308.684^{+0.003}_{-0.007}$ $<16.9$ $<6.63$
AT2018hco 58509.789 $114^{+1100}_{-58}$ $<0.474$ $<4.89$
AT2018hco 60468.668 $2073.148^{+0.003}_{-0.003}$ $<19.2$ $<198$
AT2018hco 60554.883 $2159.363^{+0.0043}_{-0.0002}$ $<29.3$ $<303$
AT2018hco 60568.988 $2173.468^{+0.006}_{-0.006}$ $<8.96$ $<92.5$
AT2018hco 60653.227 $2257.707^{+0.003}_{-0.003}$ $<13.9$ $<144$
AT2018hyz 59868.758 $1463.12^{+0.17}_{-0.17}$ $<7.58$ $<19.9$
AT2018hyz 59888.012 $1482.3^{+0.7}_{-0.7}$ $<6.1$ $<16$
AT2018hyz 60492.168 $2086.5^{+0.8}_{-0.8}$ $<7.98$ $<21$
AT2018hyz 60677.945 $2272.305^{+0.007}_{-0.007}$ $<8.9$ $<23.4$
AT2018lna 58528.910 $49^{+1100}_{-5}$ $<0.932$ $<10.3$
AT2018lna 58572.074 $91.79^{+0.04}_{-0.04}$ $<9.42$ $<104$
AT2018lna 59583.211 $1103^{+2}_{-2}$ $<3.13$ $<34.7$
AT2018lna 60439.879 $1959.594^{+0.006}_{-0.002}$ $<16.5$ $<183$
AT2018lna 60549.125 $2068.84^{+0.40}_{-0.07}$ $<3.07$ $<34$
AT2018lna 60561.590 $2081.305^{+0.003}_{-0.003}$ $<21$ $<233$
AT2019azh 59903.578 $1367.6^{+0.2}_{-0.5}$ $1.7^{+1.3}_{-0.9}$ $1.0^{+0.8}_{-0.5}$
AT2019azh 59922.633 $1386.6^{+0.1}_{-0.1}$ $<9.42$ $<5.64$
AT2019azh 59943.398 $1407.380^{+0.009}_{-0.009}$ $<8.29$ $<4.96$
AT2019azh 60462.703 $1926.685^{+0.003}_{-0.003}$ $<26.6$ $<15.9$
AT2019azh 60733.609 $2197.591^{+0.0005}_{-0.0055}$ $<22.2$ $<13.3$
AT2019dsg 59816.918 $1234.732^{+0.008}_{-0.008}$ $<6.41$ $<21.3$
AT2019dsg 60116.258 $1534.1^{+0.2}_{-0.2}$ $<6.44$ $<21.4$
AT2019dsg 60406.742 $1824.556^{+0.002}_{-2.7}$ $<11.4$ $<37.8$
AT2019dsg 60485.492 $1903.306^{+0.004}_{-0.004}$ $<42.1$ $<140$
AT2019dsg 60768.789 $2187^{+4}_{-6}$ $<0.919$ $<3.05$
AT2019dsg 60772.93 $2190.7^{+0.7}_{-0.4}$ $<1.31$ $<4.34$
AT2019ehz 60011.902 $1409.6^{+5.6}_{-0.3}$ $<2.59$ $<18.5$
AT2019eve 58744.367 $136^{+840}_{-100}$ $1.1^{+0.6}_{-0.5}$ $6^{+3}_{-2}$
AT2019teq 59867.207 $1091.09^{+0.08}_{-0.05}$ $49^{+12}_{-12}$ $505^{+121}_{-121}$
AT2019teq 59867.480 $1091.364^{+0.29}_{-0.12}$ $31^{+7}_{-7}$ $314^{+70}_{-70}$
AT2019teq 59874.238 $1098.1^{+0.7}_{-0.2}$ $18^{+4}_{-4}$ $187^{+39}_{-39}$
AT2019teq 59881.820 $1105.704^{+0.04}_{-0.09}$ $27^{+7}_{-7}$ $277^{+74}_{-74}$
AT2019teq 59881.867 $1105.751^{+0.001}_{-0.001}$ $78^{+20}_{-20}$ $1380^{+309}_{-309}$
AT2019teq 59881.871 $1105.755^{+0.002}_{-0.001}$ $82^{+15}_{-15}$ $799^{+209}_{-209}$
AT2019teq 59881.875 $1105.759^{+0.002}_{-0.003}$ $82^{+15}_{-15}$ $841^{+155}_{-155}$
AT2019teq 59887.723 $1111.6^{+0.7}_{-0.5}$ $27^{+7}_{-7}$ $273^{+72.5}_{-72.5}$
AT2019teq 59888.445 $1112.329^{+0.003}_{-0.002}$ $64^{+13}_{-13}$ $657^{+137}_{-137}$
AT2019teq 59888.895 $1112.778^{+0.001}_{-0.002}$ $71^{+14}_{-14}$ $819^{+208}_{-208}$
AT2019teq 59888.898 $1112.782^{+0.004}_{-0.002}$ $71^{+14}_{-14}$ $724^{+142}_{-142}$
AT2019teq 59895.199 $1119.083^{+0.003}_{-0.004}$ $46^{+12}_{-12}$ $343^{+90}_{-90}$
AT2019teq 59895.203 $1119.087^{+0.003}_{-0.002}$ $46^{+12}_{-12}$ $470^{+123}_{-123}$
AT2019teq 59895.207 $1119.091^{+0.002}_{-0.003}$ $52^{+13}_{-13}$ $534^{+131}_{-131}$
AT2019teq 59904.516 $1128^{+5}_{-2}$ $20^{+3}_{-3}$ $201^{+32}_{-32}$
AT2019teq 59916.148 $1140.032^{+0.003}_{-0.003}$ $62^{+13}_{-13}$ $635^{+129}_{-129}$
AT2019teq 59916.398 $1140.282^{+0.002}_{-0.002}$ $66^{+17}_{-17}$ $681^{+177}_{-177}$
AT2019teq 59916.402 $1140.286^{+0.002}_{-0.002}$ $54^{+14}_{-14}$ $553^{+144}_{-144}$
AT2019teq 59916.410 $1140.294^{+0.006}_{-0.004}$ $46^{+9}_{-9}$ $467^{+88}_{-88}$
AT2019teq 60064.922 $1288.805^{+0.002}_{-0.005}$ $48^{+11}_{-11}$ $491^{+109}_{-109}$
AT2019teq 60064.973 $1288.856^{+0.003}_{-0.003}$ $45^{+12}_{-12}$ $464^{+121}_{-121}$
AT2019teq 60064.980 $1288.864^{+0.007}_{-0.006}$ $20^{+5}_{-5}$ $203^{+51}_{-51}$
AT2019teq 60232.633 $1456.516^{+0.3}_{-0.4}$ $21^{+4}_{-4}$ $211^{+38}_{-38}$
AT2019teq 60247.367 $1471.251^{+0.003}_{-0.004}$ $33^{+9}_{-9}$ $339^{+90}_{-90}$
AT2019teq 60247.371 $1471.255^{+0.0013}_{-0.0013}$ $86^{+22}_{-22}$ $883^{+230}_{-230}$
AT2019teq 60247.375 $1471.259^{+0.004}_{-0.003}$ $38^{+9}_{-9}$ $394^{+95}_{-95}$
AT2019teq 60262.883 $1486.766^{+0.002}_{-0.003}$ $56^{+14}_{-14}$ $578^{+147}_{-147}$
AT2019teq 60262.891 $1486.774^{+0.009}_{-0.005}$ $28^{+6}_{-6}$ $283^{+62}_{-62}$
AT2019teq 60277.195 $1501.079^{+0.004}_{-0.003}$ $38^{+9}_{-9}$ $390^{+94}_{-94}$
AT2019teq 60281.414 $1505.298^{+0.004}_{-0.005}$ $24^{+7}_{-7}$ $246^{+66}_{-66}$
AT2019teq 60281.422 $1505.305^{+0.002}_{-0.003}$ $53^{+13}_{-13}$ $541^{+133}_{-133}$
AT2019teq 60467.961 $1691.8^{+1.2}_{-8.6}$ $9^{+4}_{-3}$ $90^{+40}_{-31}$
AT2019teq 60485.289 $1709.173^{+0.002}_{-0.001}$ $67^{+18}_{-18}$ $688^{+181}_{-181}$
AT2019teq 60485.293 $1709.177^{+0.0009}_{-0.0014}$ $93^{+24}_{-24}$ $959^{+250}_{-250}$
AT2019teq 60485.297 $1709.180^{+0.002}_{-0.002}$ $63^{+16}_{-16}$ $648^{+165}_{-165}$
AT2019teq 60485.301 $1709.184^{+0.002}_{-0.002}$ $78^{+17}_{-17}$ $799^{+173}_{-173}$
AT2019teq 60485.754 $1709.638^{+0.002}_{-0.003}$ $57^{+15}_{-15}$ $589^{+154}_{-154}$
AT2019teq 60485.758 $1709.641^{+0.003}_{-0.003}$ $56^{+12}_{-12}$ $575^{+121}_{-121}$
AT2020pj 58883.973 $33.4631^{+0.004}_{-10}$ $<6.67$ $<40$
AT2020pj 60464.766 $1614.256^{+0.0008}_{-0.0008}$ $<87.5$ $<525$
AT2020pj 60590.035 $1739.5^{+1.6}_{-1.6}$ $<10.5$ $<62.9$
AT2020pj 60604.438 $1753.928^{+0.0008}_{-2.3}$ $<9.26$ $<55.5$
AT2020mot 59086.973 $73^{+17}_{-12}$ $<3.16$ $<20.1$
AT2020mot 59127.027 $113^{+3}_{-3}$ $<4.38$ $<27.9$
AT2020mot 59146.957 $133^{+5}_{-5}$ $<3.13$ $<20$
AT2020mot 59250.180 $235.79^{+0.07}_{-0.07}$ $<5.17$ $<32.9$
AT2020mot 59664.453 $650.0593^{+0.005}_{-0.005}$ $<17.9$ $<114$
AT2020mot 60064.121 $1049.727^{+0.010}_{-0.010}$ $<8.89$ $<56.6$
AT2020mot 60549.977 $1535.583^{+0.002}_{-0.002}$ $<41.1$ $<262$
AT2020mot 60714.613 $1700.2^{+1.9}_{-1.9}$ $<8.22$ $<52.4$
AT2020neh 59065.730 $46^{+23}_{-34}$ $<0.364$ $<1.8$
AT2020neh 59241.113 $221.9^{+0.3}_{-13}$ $<1.66$ $<8.23$
AT2020neh 59808.578 $789.3^{+0.4}_{-0.4}$ $<4.83$ $<23.9$
AT2020neh 59823.559 $804.3^{+0.3}_{-0.3}$ $<5.31$ $<26.3$
AT2020neh 60496.129 $1476.874^{+0.003}_{-0.003}$ $<13.6$ $<67.3$
AT2020neh 60577.660 $1558.42^{+0.04}_{-0.04}$ $<6.29$ $<31.1$
AT2020neh 60701.766 $1682.51^{+0.04}_{-0.04}$ $<3.32$ $<16.4$
AT2020nov 59052.547 $25.3^{+3.4}_{-0.8}$ $3.2^{+1.0}_{-1.0}$ $30^{+9}_{-9}$
AT2020nov 59058.957 $31.7^{+0.6}_{-1.6}$ $3.8^{+1.0}_{-1.0}$ $36^{+9}_{-9}$
AT2020nov 59062.121 $34.9^{+1.5}_{-1.4}$ $3.3^{+1.0}_{-1.0}$ $31^{+9}_{-9}$
AT2020nov 59066.910 $39^{+3}_{-2}$ $3.8^{+0.9}_{-0.9}$ $36^{+8}_{-8}$
AT2020nov 59072.512 $45.3^{+1.0}_{-0.9}$ $6.0^{+1.7}_{-1.7}$ $57^{+16}_{-16}$
AT2020nov 59077.730 $50.5^{+1.6}_{-2.2}$ $3.7^{+1.2}_{-1.2}$ $35^{+11}_{-11}$
AT2020nov 59133.477 $106^{+12}_{-15}$ $1.3^{+0.6}_{-0.5}$ $13^{+6}_{-5}$
AT2020nov 59317.500 $290^{+6}_{-13}$ $6.2^{+1.7}_{-1.7}$ $58^{+16}_{-16}$
AT2020nov 59394.504 $367.27^{+0.24}_{-0.16}$ $24^{+7}_{-7}$ $228^{+62}_{-62}$
AT2020nov 59396.895 $370^{+4}_{-2}$ $21^{+4}_{-4}$ $195^{+36}_{-36}$
AT2020nov 59412.324 $385.09^{+0.02}_{-0.06}$ $17^{+4}_{-4}$ $162^{+35}_{-35}$
AT2020nov 59416.176 $388.9^{+2.4}_{-1.0}$ $16^{+3}_{-3}$ $150^{+27}_{-27}$
AT2020nov 59421.953 $394.721^{+0.004}_{-0.004}$ $29^{+8}_{-8}$ $267^{+72}_{-72}$
AT2020nov 59496.199 $468.97^{+0.04}_{-0.04}$ $10^{+3}_{-3}$ $94^{+25}_{-25}$
AT2020nov 59496.316 $469.08^{+0.57}_{-0.03}$ $21^{+5}_{-5}$ $193^{+43}_{-43}$
AT2020nov 59783.355 $756.1229^{+0.009}_{-0.009}$ $7^{+3}_{-3}$ $67^{+29}_{-24}$
AT2020nov 59797.652 $770.42^{+0.17}_{-0.11}$ $15^{+3}_{-3}$ $137^{31}_{-31}$
AT2020nov 59799.625 $772.4^{+4.1}_{-1.4}$ $9.6^{+1.9}_{-1.9}$ $90^{+18}_{-18}$
AT2020nov 59817.082 $789.8^{+0.7}_{-3.8}$ $4.5^{+1.5}_{-1.5}$ $42^{+14}_{-14}$
AT2020nov 59830.562 $803^{+15}_{-6}$ $3.6^{+1.4}_{-1.1}$ $34^{+13}_{-11}$
AT2020nov 60066.664 $1039.4^{+4}_{-0.4}$ $5^{+3}_{-2}$ $45^{+27}_{-20}$
AT2020nov 60697.340 $1670^{+3}_{-7}$ $5.4^{+1.9}_{-1.6}$ $50^{+18}_{-15}$
AT2020nov 60464.645 $1437.412^{+0.007}_{-0.005}$ $<15.1$ $<141$
AT2020nov 60482.066 $1454.834^{+0.003}_{-0.004}$ $<22.7$ $<213$
AT2020opy 59094.102 $56^{+47}_{-24}$ $0.6^{+0.2}_{-0.2}$ $21^{+7}_{-7}$
AT2020opy 60121.434 $1083.201^{+0.010}_{-0.010}$ $<5.18$ $<191$
AT2020opy 60460.504 $1422.272^{+0.003}_{-0.003}$ $<22.9$ $<845$
AT2020opy 60598.016 $1559.8^{+1.9}_{-1.7}$ $<4.63$ $<171$
AT2020wey 59163.875 $33^{+3}_{-14}$ $<1.29$ $<1.18$
AT2020wey 59191.008 $61^{+11}_{-11}$ $<1.63$ $<1.49$
AT2020wey 60403.840 $1273.35^{+0.04}_{-0.04}$ $<9.91$ $<9.09$
AT2020wey 60542.402 $1411^{+3}_{-3}$ $<2.2$ $<2.02$
AT2020wey 60742.219 $1612^{+3}_{-4}$ $<4.19$ $<3.84$
AT2020wey 60751.469 $1621^{+0.5}_{-0.5}$ $<6.52$ $<5.98$
AT2020vwl 59235.754 $103.026^{+0.011}_{-15}$ $<1.97$ $<2.99$
AT2020vwl 59248.641 $115.9^{+1.2}_{-1.2}$ $<2.54$ $<3.86$
AT2020vwl 59324.539 $192^{+14}_{-22}$ $<1.38$ $<2.1$
AT2020vwl 59402.941 $270^{+3}_{-3}$ $<5.0$ $<7.59$
AT2020vwl 59420.695 $287.97^{+0.07}_{-0.07}$ $<10.6$ $<16.1$
AT2020vwl 59589.926 $457.198^{+0.010}_{-0.010}$ $<8.28$ $<12.6$
AT2020vwl 59831.680 $699^{+4}_{-4}$ $<2.08$ $<3.15$
AT2020vwl 59849.492 $717^{+2}_{-2}$ $<2.56$ $<3.89$
AT2020vwl 59865.191 $732^{+3}_{-3}$ $<2.6$ $<3.94$
AT2020vwl 59904.930 $772.20^{+0.07}_{-0.07}$ $<7.51$ $<11.4$
AT2020vwl 59918.219 $785.49^{+0.04}_{-0.04}$ $<4.47$ $<6.78$
AT2020vwl 60122.688 $989.960^{+0.007}_{-0.007}$ $<6.64$ $<10.1$
AT2020vwl 60573.629 $1441^{+5}_{-5}$ $<9.8$ $<14.9$
AT2020vwl 60588.379 $1456^{+4}_{-4}$ $<1.67$ $<2.54$
Appendix C MOSFit Parameters
The MOSFiT TDE model has nine free physical parameters (Mockler et al., 2019; Nicholl et al., 2022), of which eight are of potential interest for discriminating among competing TDE emission models: the viscous delay time $t_{v}$, the disruption impact parameter $b$, the SMBH mass $M_{\rm BH,mosfit}$, the column density along the line of sight in the host galaxy $n_{\rm H,host}$, the normalization of the photospheric radius $R_{\rm ph,0}$ at the time when the observed bolometric luminosity of the TDE reaches $L_{\rm Edd}$ and the corresponding power law exponent $l_{\rm ph}$ that relates the photospheric radius to the instantaneous luminosity as $R_{\rm ph}\propto R_{\rm ph,0}(L/L_{\rm Edd})^{l_{\rm ph}}$, the mass of the disrupted star $M_{*}$, and the efficiency with which accreting mass is converted to radiation, $\epsilon$. The remaining parameter is the time of first fallback relative to the first data point in the light curve, $t_{0}$. Although not directly physically relevant, $t_{0}$ can be used in combination with the light curves to calculate the date of disruption, MJD0, and the number of rest-frame days between disruption and peak optical/UV light, $t_{\rm rise}$ (which has been suggested to correlate with other physical properties of TDEs; e.g., Rees 1988; van Velzen et al. 2021). We derive two additional parameters of possible physical interest from our MOSFiT solutions: $\dot{M}_{\rm max}$, the maximum accretion rate onto the SMBH during each TDE, and $L_{\rm max}$, the maximum bolometric luminosity during the TDE. For our analysis, we normalize $\dot{M}_{\rm max}$ by $\dot{M}_{\rm Edd}$ (as defined in Section 3.3) to remove the dependence on the efficiency.
The best-fit physically relevant parameters and $1\sigma$ confidence intervals for our fits are given in Table 7. We note that our results for the TDEs also studied by Nicholl et al. (2022) are similar to those reported in that work. For completeness, we also include a table of the results for all of the Anderson-Darling tests conducted for our dataset as described in Section 3.2.
[…]
=1.8in {rotatetable}
Table 7: MOSFiT Parameters
TDE Name $t_{v}$ (d) $b$ $\log(M_{\rm BH,mosfit}/$ $\log{(n_{\rm H,host}/}$ $R_{\rm ph,0}$ $M_{*}$ $\epsilon$ $t_{\rm rise}$ (d) $\dot{M}_{\rm max}/$ $L_{\rm max}$
(rest-frame) $M_{\odot})$ cm-2) ($M_{\odot}$) (rest-frame) $\dot{M}_{\rm Edd}$ ($10^{44}$ erg s-1)
ASASSN-14ae $0.14^{+2.35}_{-0.13}$ $0.30^{+0.12}_{-0.10}$ $6.0^{+0.3}_{-0.2}$ $18.1^{+1.5}_{-1.3}$ $4.8^{+0.8}_{-0.7}$ $1.0^{+0.2}_{-0.2}$ $0.009^{+0.019}_{-0.005}$ $27^{+9}_{-11}$ $0.7^{+1.0}_{-0.4}$ $0.65^{+0.07}_{-0.13}$
ASASSN-14li $0.12^{+6.48}_{-0.12}$ $0.64^{+0.21}_{-0.08}$ $7.62^{+0.05}_{-0.11}$ $21.11^{+0.02}_{-0.02}$ $0.7^{+0.6}_{-0.3}$ $0.8^{+0.2}_{-0.2}$ $0.23^{+0.12}_{-0.13}$ $132^{+10}_{-16}$ $0.5^{+0.3}_{-0.2}$ $21^{+9}_{-6}$
iPTF15af $0.4^{+18.3}_{-0.4}$ $0.7^{+0.6}_{-0.3}$ $7.1^{+0.2}_{-0.3}$ $21.00^{+0.05}_{-0.06}$ $0.20^{+0.18}_{-0.07}$ $1.4^{+0.9}_{-0.8}$ $0.07^{+0.16}_{-0.06}$ $28^{+12}_{-16}$ $1.3^{+3.8}_{-1.1}$ $8^{+5}_{-7}$
ASASSN-15oi $0.05^{+1.09}_{-0.05}$ $0.94^{+0.02}_{-0.06}$ $7.22^{+0.09}_{-0.11}$ $17.6^{+1.1}_{-0.9}$ $83^{+131}_{-45}$ $1.00^{+0.05}_{-0.05}$ $0.031^{+0.006}_{-0.006}$ $131^{+15}_{-8}$ $0.6^{+0.2}_{-0.2}$ $8.5^{+0.8}_{-1.1}$
iPTF16axa $0.2^{+3.2}_{-0.2}$ $0.93^{+0.11}_{-0.13}$ $7.28^{+0.13}_{-0.12}$ $21.17^{+0.02}_{-0.03}$ $1.0^{+0.9}_{-0.4}$ $0.5^{+0.4}_{-0.2}$ $0.24^{+0.11}_{-0.10}$ $142^{+29}_{-31}$ $1.3^{+1.0}_{-0.6}$ $15^{+4}_{-5}$
PS16dtm $84^{+26}_{-25}$ $0.63^{+0.14}_{-0.19}$ $7.3^{+0.2}_{-0.2}$ $17.3^{+1.3}_{-0.8}$ $4^{+2}_{-1}$ $0.5^{+0.3}_{-0.1}$ $0.05^{+0.05}_{-0.02}$ $104^{+7}_{-11}$ $0.2^{+0.2}_{-0.1}$ $2.51^{+0.08}_{-0.08}$
iPTF16fnl $0.07^{+0.51}_{-0.06}$ $1.04^{+0.05}_{-0.57}$ $6.1^{+0.3}_{-0.4}$ $20.8^{+0.1}_{-3.0}$ $3.1^{+1.5}_{-0.8}$ $0.105^{+1.113}_{-0.008}$ $0.005^{+0.003}_{-0.001}$ $17^{+4}_{-9}$ $0.4^{+0.8}_{-0.1}$ $0.5^{+0.3}_{-0.6}$
OGLE17aaj $5^{+25}_{-5}$ $0.98^{+0.12}_{-0.16}$ $7.22^{+0.13}_{-0.29}$ $19^{+2}_{-1}$ $0.13^{+0.10}_{-0.03}$ $0.7^{+1.1}_{-0.4}$ $0.03^{+0.07}_{-0.02}$ $14^{+3}_{-4}$ $0.5^{+0.9}_{-0.3}$ $6.4^{+3.2}_{-5.0}$
AT2017eqx $0.04^{+0.91}_{-0.04}$ $1.0^{+0.1}_{-0.2}$ $6.7^{+0.3}_{-0.3}$ $21.0^{+0.1}_{-2.9}$ $5^{+26}_{-3}$ $0.2^{+0.3}_{-0.1}$ $0.02^{+0.05}_{-0.01}$ $39^{+13}_{-17}$ $0.3^{+0.8}_{-0.2}$ $1.9^{+1.5}_{-2.6}$
AT2018zr $0.01^{+0.31}_{-0.01}$ $0.87^{+0.07}_{-0.07}$ $6.79^{+0.05}_{-0.11}$ $20.92^{+0.05}_{-0.07}$ $9^{+3}_{-2}$ $0.11^{+0.07}_{-0.04}$ $0.04^{+0.02}_{-0.02}$ $8.0^{+0.7}_{-0.8}$ $0.2^{+0.1}_{-0.0}$ $1.5^{+0.2}_{-0.3}$
AT2018bsi $0.1^{+3.1}_{-0.1}$ $0.9^{+0.3}_{-0.2}$ $6.9^{+0.3}_{-0.4}$ $20.6^{+0.2}_{-0.8}$ $1.5^{+2.2}_{-0.9}$ $0.7^{+0.6}_{-0.2}$ $0.011^{+0.033}_{-0.005}$ $78^{+32}_{-58}$ $0.3^{+0.8}_{-0.2}$ $2.1^{+1.0}_{-2.3}$
AT2018dyb $0.06^{+0.53}_{-0.06}$ $0.90^{+0.08}_{-0.03}$ $7.15^{+0.04}_{-0.04}$ $17.2^{+0.7}_{-0.7}$ $1.9^{+0.2}_{-0.2}$ $0.33^{+0.04}_{-0.09}$ $0.23^{+0.10}_{-0.04}$ $25^{+2}_{-4}$ $1.1^{+0.2}_{-0.2}$ $10.8^{+0.6}_{-0.7}$
AT2018hco $0.1^{+2.4}_{-0.1}$ $0.77^{+0.10}_{-0.10}$ $6.87^{+0.08}_{-0.27}$ $19^{+2}_{-2}$ $1.7^{+0.2}_{-0.2}$ $0.12^{+0.10}_{-0.05}$ $0.08^{+0.05}_{-0.04}$ $8^{+1}_{-2}$ $0.32^{+0.27}_{-0.09}$ $2.4^{+0.3}_{-0.5}$
AT2018hyz $0.10^{+1.13}_{-0.09}$ $0.75^{+0.04}_{-0.23}$ $6.21^{+0.43}_{-0.08}$ $19^{+1}_{-2}$ $4.6^{+0.8}_{-0.4}$ $1.2^{+0.1}_{-0.8}$ $0.0035^{+0.0089}_{-0.0007}$ $7^{+2}_{-3}$ $1.6^{+0.6}_{-1.4}$ $1.31^{+0.07}_{-0.09}$
AT2018lna $0.10^{+1.90}_{-0.09}$ $0.97^{+0.12}_{-0.19}$ $6.65^{+0.15}_{-0.15}$ $18^{+2}_{-2}$ $2.0^{+1.1}_{-0.5}$ $0.6^{+0.3}_{-0.2}$ $0.007^{+0.003}_{-0.003}$ $15^{+5}_{-7}$ $0.4^{+0.3}_{-0.1}$ $1.7^{+0.2}_{-0.4}$
ASASSN-19bt $1.6^{+1.2}_{-0.6}$ $0.89^{+0.12}_{-0.14}$ $6.80^{+0.05}_{-0.11}$ $20.79^{+0.03}_{-0.10}$ $8^{+3}_{-4}$ $0.7^{+1.0}_{-0.1}$ $0.009^{+0.004}_{-0.005}$ $9.5^{+0.9}_{-1.2}$ $0.4^{+0.2}_{-0.1}$ $2.5^{+0.4}_{-0.2}$
AT2019azh $0.06^{+0.63}_{-0.06}$ $0.32^{+0.24}_{-0.05}$ $6.77^{+0.09}_{-0.06}$ $19.2^{+0.9}_{-2.1}$ $1.0^{+0.2}_{-0.1}$ $0.6^{+0.2}_{-0.2}$ $0.13^{+0.09}_{-0.04}$ $9^{+1}_{-2}$ $0.9^{+0.5}_{-0.2}$ $3.7^{+0.3}_{-0.9}$
AT2019dsg $0.1^{+2.5}_{-0.1}$ $0.89^{+0.18}_{-0.13}$ $7.07^{+0.14}_{-0.17}$ $19^{+1}_{-2}$ $0.6^{+0.2}_{-0.1}$ $0.5^{+0.3}_{-0.2}$ $0.08^{+0.13}_{-0.05}$ $36^{+10}_{-12}$ $0.9^{+0.7}_{-0.3}$ $7.7^{+1.7}_{-2.4}$
AT2019ehz $0.2^{+1.3}_{-0.2}$ $0.46^{+0.10}_{-0.07}$ $6.69^{+0.04}_{-0.04}$ $20.92^{+0.04}_{-0.05}$ $2.7^{+0.5}_{-0.4}$ $0.14^{+0.02}_{-0.03}$ $0.28^{+0.07}_{-0.09}$ $10^{+1}_{-2}$ $1.1^{+0.5}_{-0.4}$ $3.6^{+0.7}_{-0.5}$
AT2019eve $0.04^{+0.67}_{-0.03}$ $0.14^{+0.13}_{-0.06}$ $6.4^{+0.2}_{-0.2}$ $19^{+2}_{-2}$ $10^{+24}_{-5}$ $0.5^{+0.4}_{-0.2}$ $0.02^{+0.04}_{-0.02}$ $16^{+4}_{-4}$ $0.03^{+0.03}_{-0.01}$ $0.10^{+0.01}_{-0.04}$
AT2019qiz $5^{+1}_{-4}$ $0.40^{+0.07}_{-0.26}$ $6.19^{+0.07}_{-0.16}$ $20.6^{+0.2}_{-0.7}$ $3.6^{+2.9}_{-0.5}$ $0.9^{+2.6}_{-0.2}$ $0.004^{+0.002}_{-0.002}$ $7^{+1}_{-3}$ $0.22^{+0.09}_{-0.08}$ $0.3^{+0.1}_{-0.2}$
AT2019teq $16^{+13}_{-16}$ $1.1^{+0.4}_{-0.3}$ $5.8^{+0.3}_{-0.2}$ $20.93^{+0.13}_{-0.12}$ $4^{+3}_{-2}$ $0.7^{+0.6}_{-0.4}$ $0.006^{+0.012}_{-0.004}$ $21^{+8}_{-18}$ $5^{+8}_{-3}$ $0.6^{+0.1}_{-0.5}$
AT2020pj $0.07^{+0.94}_{-0.06}$ $0.8^{+0.2}_{-0.2}$ $6.5^{+0.2}_{-0.2}$ $21.21^{+0.06}_{-0.07}$ $5^{+7}_{-2}$ $0.4^{+0.2}_{-0.2}$ $0.010^{+0.014}_{-0.006}$ $17^{+5}_{-9}$ $0.5^{+1.3}_{-0.3}$ $1.4^{+0.6}_{-0.7}$
AT2020mot $15^{+2}_{-3}$ $0.13^{+0.04}_{-0.04}$ $6.81^{+0.05}_{-0.06}$ $20.76^{+0.07}_{-0.12}$ $0.78^{+0.10}_{-0.06}$ $11^{+4}_{-4}$ $0.05^{+0.04}_{-0.03}$ $20^{+2}_{-2}$ $0.4^{+0.1}_{-0.1}$ $2.1^{+0.5}_{-0.4}$
AT2020neh $0.05^{+0.40}_{-0.05}$ $0.3^{+0.4}_{-0.1}$ $6.18^{+0.14}_{-0.14}$ $18^{+2}_{-1}$ $10^{+300}_{-7}$ $0.6^{+0.4}_{-0.2}$ $0.02^{+0.10}_{-0.02}$ $8^{+3}_{-5}$ $0.8^{+3.7}_{-0.7}$ $0.6^{+0.4}_{-1.2}$
AT2020nov $1.0^{+3.1}_{-0.8}$ $0.99^{+0.04}_{-0.08}$ $7.26^{+0.03}_{-0.07}$ $18^{+2}_{-2}$ $3.3^{+0.6}_{-0.9}$ $0.31^{+0.13}_{-0.04}$ $0.010^{+0.002}_{-0.002}$ $16^{+2}_{-2}$ $0.034^{+0.050}_{-0.004}$ $0.85^{+0.05}_{-0.58}$
AT2020opy $0.4^{+3.6}_{-0.4}$ $0.97^{+0.10}_{-0.10}$ $6.58^{+0.13}_{-0.15}$ $18.3^{+0.9}_{-1.3}$ $2.2^{+0.4}_{-0.2}$ $1.0^{+0.4}_{-0.2}$ $0.0038^{+0.0022}_{-0.0008}$ $16^{+4}_{-8}$ $0.6^{+0.4}_{-0.2}$ $2.0^{+0.1}_{-0.1}$
AT2020wey $0.1^{+6.1}_{-0.1}$ $0.8^{+0.2}_{-0.2}$ $6.59^{+0.11}_{-0.14}$ $20.74^{+0.08}_{-0.14}$ $24^{+84}_{-19}$ $0.6^{+0.3}_{-0.1}$ $0.006^{+0.003}_{-0.002}$ $19^{+3}_{-3}$ $0.4^{+0.5}_{-0.2}$ $1.4^{+0.5}_{-0.6}$
AT2020vwl $0.1^{+2.1}_{-0.1}$ $0.18^{+0.05}_{-0.05}$ $6.7^{+0.2}_{-0.2}$ $20.1^{+0.7}_{-3.1}$ $3^{+4}_{-1}$ $1.1^{+1.8}_{-0.3}$ $0.03^{+0.04}_{-0.02}$ $90^{+25}_{-29}$ $0.07^{+0.20}_{-0.04}$ $0.4^{+0.1}_{-0.7}$
Table 8: Anderson-Darling Test Results. The numbers give the probabilities that the underlying parameter distributions are the same for the two sub-samples being compared in each column. Entries for which $p\leq 0.05$ are in bold.
Early vs. Late No Detection + Host vs. Late No Detection + Host vs. Early
$M_{\rm BH,host}$ 0.050 0.049 0.227
$\log M_{\rm BH,mosfit}$ $>0.25$ 0.199 $>0.25$
$t_{\rm viscous}$ 0.188 $>0.25$ $>0.25$
$b$ $>0.25$ 0.196 $>0.25$
$\log n_{\rm H,host}$ 0.132 0.074 $>0.25$
$R_{\rm ph,0}$ 0.026 0.056 $>0.25$
$M_{*}$ $>0.25$ $>0.25$ $>0.25$
$\epsilon$ $>0.25$ $>0.25$ $>0.25$
$l_{\rm photo}$ $>0.25$ $>0.25$ $>0.25$
$t_{\rm rise}$ $>0.25$ 0.151 $>0.25$
$\dot{M}_{\rm max}$ $>0.25$ $>0.25$ $>0.25$
$L_{\rm max}$ $>0.25$ $>0.25$ $>0.25$